### v3

# Compact moduli of Calabi-Yau cones and Sasaki-Einstein spaces

Abstract.
We construct proper moduli algebraic spaces of K-polystable $\mathbb{Q}$-Fano cones (a.k.a. Calabi-Yau cones) or equivalently their links i.e., Sasaki-Einstein manifolds with singularities.
As a byproduct, it gives alternative algebraic construction of proper K-moduli of $\mathbb{Q}$-Fano varieties. In contrast to the previous algebraic proof of its properness ([BHLLX21, LXZ22]), we do not use the $\delta$-invariants ([FO18, BJ22]) nor the $L^{2}$-normalized Donaldson-Futaki invariants. We use the local normalized volume of [Li18] and the higher $\Theta$-stable reduction [Od24b] instead.
2000 Mathematics Subject Classification:
Primary 14D20; Secondary 53C25.
1. Introduction
It follows from the recent breakthrough [DS14], combined with the Gromov’s precompactness theorem and the theory of Gromov-Hausdorff distance [Gro99], that there should be compactifying topological space of moduli spaces of Kähler-Einstein Fano manifolds, letting the boundary parametrizes certain singular Kähler-Einstein Fano varieties. Indeed, [DS14] uses Cheeger-Colding theory to prove that for Kähler-Einstein smooth Fano manifolds of fixed dimensions, the Gromov-Hausdorff limits, which always exists for subsequences, admit the structure of log terminal Kähler-Einstein Fano varieties, which are $\mathbb{Q}$-Gorenstein smoothable (cf., [OSS16, §2]). In particular, they admit the structure of algebraic varieties. This also matches the general K-moduli conjecture then ([Od10, Conjecture 5.2]), a purely algebro-geometric conjecture, which was later refined in the course of developments. Indeed, the obtained compact moduli topological spaces were enhanced to be proper (compact) algebraic spaces by using K-stability in the series of works [OSS16, LWX19, SSY16, Od15a] (cf., also [MM93] which predates K-stability). By more birational algebro-geometric discussions after that, recently we witnessed a celebrated algebraic (re-)proof and generalization of the facts that the moduli (stack) of K-polystable Fano varieties satisfies the valuative criterion of properness by [BHLLX21, LXZ22], and the (categorical) good moduli 111categorical moduli of an algebraic stack means the one in e.g., [Stacksproject, 106.12] and good moduli is in the sense of [Alp13]. The author confesses that he has often used the term “coarse moduli” before, for the exactly same meaning as the categorical moduli in older papers, i.e., not assuming the bijectivity condition for geometric points, which is assumed in most of the literature. Henceforth, we will avoid terminlogical confusion. is projective ([CP21, XZ22]). Their algebraic proof depends on the theory of $\delta$-invariants [FO18, BJ22] of Fano varieties. This reproves and extends the original construction [LWX19, Od15a] (cf., also closely related results in [SSY16]) to generally singular $\mathbb{Q}$-Fano varieties.
In this paper, we extend this K-moduli construction to that of Fano cones, e.g., contracted (pluri-)anticanonical line bundles over Kähler-Einstein Fano manifolds, by a different method. What underlies historically behind this theory of cone type varieties is the real odd-dimensional version of Kähler-Einstein manifolds, the structure of Sasaki-Einstein manifolds on their links (see the textbook [BG07]) as we review in the second section. One benefit of our approach is that it also gives alternative proof of the algebraic construction of the K-moduli of Fano varieties, without depending on the $\delta$-invariant nor $L^{2}$-normalized Donaldson-Futaki invariants, but rather we only use the normalized volume or the volume density of the Einstein metrics, to match more easily with theory of log terminal singularities. We work over an algebraically closed field $\mathbbm{k}$ of characteristic 00.
Main Theorem (cf., Definitions 2.9, 2.15, 2.21, Theorem 3.14).
For each algebraically closed field $\mathbbm{k}$ of characteristic 00, $n\in\mathbb{Z}_{>0}$ and $V\in\mathbb{R}_{>0}$, there is a proper moduli algebraic $\mathbbm{k}$-space of $n$-dimensional K-polystable Fano cones over $\mathbbm{k}$ (a.k.a., Calabi-Yau cones222 in some literature, it requires slightly more constraints for this terminology e.g., [CH22]) of the volume density $V$.
If $\mathbbm{k}=\mathbb{C}$, equivalently, real $(2n-1)$-dimensional compact Sasaki-Einstein spaces of volume density $V$ form a compact Moishezon analytic space for each fixed $n$ and $V$.
More precise version, logarithmic generalization and a corollary are stated and proved as Theorem 3.14, 3.17 and 3.18 respectively. We conjecture (Conjecture 3.20) that these good moduli spaces are also projective schemes with Weil-Petersson type (singular) Kähler metrics.
2. Preparation
We give a brief review of some preparatory materials. Most of (but perhaps not all of) the materials in this section are known before as we give references to each.
2.1. Sasakian geometry
A classical differential geometric (or contact geometric) counterpart of the Ricci-flat Kähler cone is the geometry of Sasakian manifolds, which we briefly recall. This subsection is intended to be more differential geometric review, and only works over $\mathbbm{k}=\mathbb{C}$ whenever some varieties appear.
Definition 2.1 (Sasakian manifolds cf., [Sas60, BG14]).
A Riemannian manifold $(S,g)$ of (real) odd dimension $12n-$ 1 with $n\in\mathbb{Z}_{>0}$ is said to underlie a Sasakian manifold if there is a complex structure $J$ on $C^{o}(S):=S\times\mathbb{R}_{>0}$ with respect to which $g_{C(S)}$ is a Kähler metric, so that $J$ further extends to the cone $C(S):=C^{o}(S)\cup 0$. The corresponding Kähler form is, by the definition of $g_{C(S)}$, $\omega_{C(S)}:=\sqrt{-1}\partial\overline{\partial}r^{2}$.
Then, the corresponding $\xi:=J\bigl{(}r\frac{\partial}{\partial r}\bigr{)}$ is called the Reeb vector field. To avoid confusion with more algebraic flexible variant (positive vector field in Definition 2.9), we call it the metric Reeb vector field. We also have a contact $1$-form $\eta:=\iota_{\xi}g=\iota_{\frac{\partial}{\partial r}}\omega_{C(S)}$ associated, where $\iota$ stands for the interior product. The actual Sasaki structure on $M$ is in addition to $g$, further we encode $\xi$ (or $\eta$ equivalently). 333although, in most of quasi-regular cases, this is redundant by [BGN03, 8.4], [BG05, 20, 21]
There are also associated structures such as $(1,1)$ type real tensor field $\Phi\in\Gamma(M,{\rm End}(T_{S}))$, which satisfies $\Phi\circ\Phi=-{\rm id}+\xi\otimes\eta$, where ${\rm End}(T_{S})$ means the endomorphism bundle of the (real) tangent bundle, ${\rm id}$ means the identity map and the contact form $\eta$ is defined as $\eta(v)=g(\xi,v)$. The datum is again equivalent to the (almost) complex structure $J$ on $C(S)$ by the formula $\Phi(v)=J(v-g(\xi,v)\xi)=\nabla^{\rm LC}_{v}\xi$ by the Levi-Civita connection $\nabla^{\rm LC}$. Below, we assume $S$ is compact throughout, unless otherwise stated.
Definition 2.2 (Quasi-regular, irregular).
$\overline{\mathbb{R}\xi}\subset{\rm Isom}(S,g)$ is known to be a $r(\xi)$-dimensional compact torus. We denote its complexification as $T(\mathbb{C})\simeq(\mathbb{C}^{*})^{r(\xi)}$. If $r(\xi)=1$ (resp., $r(\xi)>1$), we call $(S,g)$ or corresponding Fano cone is quasi-regular (resp., irregular). If $(S,g)$ is quasi-regular and further the action of $T(\mathbb{C})\simeq\mathbb{C}^{*}$ on $C(S)$ is free, we call $(S,g)$ is regular.
The following is also well-known (cf., e.g., [BG07]).
Proposition 2.3.
For a compact Sasakian manifold $(S,g,\xi)$, the following are equivalent:
  1. ($\mathrm{i}$)
    $(C(S),g_{C(S)})$ is Ricci-flat.
  2. ($\mathrm{ii}$)
    $(S,g)$ satisfies $\operatorname{Ric}_{g}=(\operatorname{dim}(X)-1)g$, hence Einstein manifold in particular (called Sasaki-Einstein manifold).
  3. ($\mathrm{iii}$)
    In the quasi-regular case, $S/(T(\mathbb{C})\simeq\mathbb{C}^{*})$ admits a natural branch divisor with standard coefficients $\Delta$ (which is 00 in the regular case) so that $(S/(T(\mathbb{C})\simeq\mathbb{C}^{*}),\Delta)$ is a log K-polystable $\mathbb{Q}$-Fano variety with a conical weak Kähler-Einstein metric.
In the above case, $C(S)\cup 0$ has only log terminal singularity at the vertex.
As our highlight, we consider the limiting behaviour.
Definition 2.4.
We consider all the (real) $12n-$ 1-dimensional compact Sasaki-Einstein manifolds $(S,g)$ of the volume $V$ and denote their isomorphic classes as $M_{n,V}$.
Corollary 2.5.
$M_{n,V}$ is precompact with respect to the Gromov-Hausdorff topology. Furthermore, $M_{n,V}(i=1,2,\cdots)$ satisfies the non-collapsing condition (compare [DS14]):
For any $0<r\leq\operatorname{diam}(S_{i},g_{i})$, there is a positive real number $c$ such that we have $\operatorname{vol}(B_{r}(p,(S_{i},g_{i}))\geq cr^{n}$. Here, $\operatorname{diam}(-)$ means the diameter, $\operatorname{vol}$ means the volume and $(B_{r}(p,(S_{i},g_{i}))$ means the geodesic ball of radius $r$ and the center $p$ in $(S_{i},g_{i})$.
Proof.
The former claim follows from the Myers theorem and Gromov precompactness theorem. The latter claim follows from the Bishop-Gromov comparison theorem. ∎
From this claim, we naturally expect some compactness theorem and the main aim of our paper is to give a proof to its algebraic analogue.
Martelli-Sparks-Yau [MSY08] studied necessary condition of the existence of Sasaki-Einstein metric (using the case study when $(C(S),J)$ in [MSY06]) and showed that the volume is an algebraic number.
Definition 2.6.
For a compact Sasaki-Einstein manifold $(S,g,\xi)$, define its volume as
$\operatorname{vol}^{\rm DG}(\xi)$ $:=\frac{1}{(2\pi)^{n}n!}\int_{C(S)}e^{-r^{2}}\omega_{C(S)}^{n}$
$=\frac{\operatorname{vol}(S,g)}{\operatorname{vol}{(S^{2n-1}(1))}}.$
The denominator is with respect to the standard round metric on the unit sphere. Note that the above invariant ${\rm vol}^{\rm DG}$ is invariant under the rescale of $r$ by $cr(c\in\mathbb{R}_{>0})$.
The above volume has a relatively algebraic nature and is determined only by a positive vector field $\xi$ as follows. We leave the details of the proof of the following to [CS18, §6], [MSY08] or [Li18].
Proposition 2.7 ([MSY08],[CS18], [CS19, Proposition 6.6], and Lemma 2.14 later).
For any compact Sasaki manifold $(S,g,\xi)$, the cone $(C(S),J)$ is a complex affine variety, which we write as $\mathop{\mathrm{Spec}}\nolimits(R)$, acted by the algebraic $\mathbbm{k}$-torus $T(\mathbb{C})$ which induces the $T$-eigenspaces decomposition $R=\oplus_{\vec{m}\in{\rm Hom}(T(\mathbb{C}),\mathbb{C}^{*})}$ and we regard $\xi$ as an element of ${\rm Hom}(\mathbb{C}^{*},T(\mathbb{C}))$. Then, we can define and write the index character $F$ ([MSY08]) as
$F(\xi,t)$ $:=\sum_{\vec{m}\in M}e^{-t\langle\vec{m},\xi\rangle}\operatorname{dim}R_{\vec{m}}$
$=\frac{A_{0}(\xi)}{t^{n}}-\frac{A_{1}(\xi)}{t^{n-1}}+O(t^{2-n})$
and further, if $\xi$ is the metric Reeb vector field for the Ricci-flat Kähler metric, we have
(1) $A_{0}(\xi)$ $=\operatorname{vol}^{\rm DG}(\xi)$
(2) $=\widehat{\operatorname{vol}}({\rm val}_{\xi}).$
Here (2) is in the sense of C.Li [Li18], which algebraizes the theory and is the topic of the next subsection. Actually the equality $A_{0}(\xi)=\widehat{\operatorname{vol}}({\rm val}_{\xi})$ makes sense and holds true for more general cases (abstract Reeb vector field in Definition 2.9) as we see in Lemma 2.14.
In particular the above volume $\operatorname{vol}^{\rm DG}(\xi)$ is determined only by the multi-Hilbert function of the decomposition $\Gamma(\mathcal{O}_{X})=R=\oplus_{\vec{m}}R_{\vec{m}}$.
Theorem 2.8.
If $(S,g)$ is a compact Sasaki-Einstein manifold, the following hold.
  1. ($\mathrm{i}$)
    ([MSY06, MSY08]) $\operatorname{vol}^{\rm DG}(-)$ extends to whole Reeb cone (cf., Definition 2.9) and is minimized at $\xi$ which is determined by the Sasakian structure.
  2. ($\mathrm{ii}$)
    ([Li18, CS19]) the metric Reeb vector field $\xi$ of the metric tangent cone $C(S)$ satisfies the normalization (or gauge-fixing) condition $A_{C(S)}(\xi)=n$. Here, $A_{C(S)}(-)$ denotes the log discrepancy function in the theory of the minimal model program and ${\rm val}_{\xi}$ denotes the naturally corresponding valuation of $\mathcal{O}_{C(S),0}$ to $\xi$.
The above results show some algebro-geometric natures of the Sasakian geometry. From the next section, we mainly work by such algebro-geometric framework.
2.2. Affine cone
Henceforth, we work on more algebro-geometric side, over an algebraically closed field $\mathbbm{k}$ of characteristic 00 unless otherwise stated. Take an arbitrary algebraic torus $T$, we set $N:={\rm Hom}(\mathbb{G}_{m},T)$, $M:=N^{\vee}={\rm Hom}(T,\mathbb{G}_{m})$ throughout. Before going further, we promise the further toric notation throughout the paper (which follows [Od24b]).
Notations 1.
For any rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$, we denote an affine toric variety corresponding to it as $U_{\tau}(\supset T)$ and its $T$-invariant vertex (closed point) as $p_{\tau}$. We denote the quotient stack $[U_{\tau}/T]$ as $\Theta_{\tau}$. We often consider non-zero irrational element $\xi\in(\tau\setminus N_{\mathbb{Q}}:=N\otimes_{\mathbb{Z}}\mathbb{Q})$ i.e., of $\mathbb{Q}$-rank $r$.
On the dual side, we set $M:={\rm Hom}(N,\mathbb{Z})$ as the dual lattice, $\tau^{\vee}:=\{x\in M\otimes\mathbb{R}\mid\langle\tau,x\rangle\subset\mathbb{R}_{\geq 0}\}$, $\mathcal{S}_{\tau}:=\tau^{\vee}\cap M$. If we regard $M$ as the the character lattice of $T$, then the character of $T$ which corresponds to $\vec{m}\in M$ is denoted by $\chi_{\vec{m}}$.
Definition 2.9 (cf., [LS13, CS18, CS19]).
  1. ($\mathrm{i}$)
    Consider a normal affine variety $X$ with the vertex $x$ and an action of an algebraic torus $T$, which is good in the sense of [LS13], i.e., the action is effective and $x$ is contained in the closure of any $T$-orbit which characterizes $x$.  [^4] When $r=1$, $T\curvearrowright X\ni x$ is also sometimes said to be a quasicone in the literature. We consider the decomposition
    $R:=\Gamma(X,\mathcal{O}_{X})=\oplus_{\vec{m}\in M}R_{\vec{m}}$
    and define the moment monoid as
    $\sigma_{R}:=\{\vec{m}\in M\mid R_{\vec{m}}\neq 0\},$
    the moment cone as
    $\mathbb{R}_{\geq 0}\sigma_{R}\subset M\otimes\mathbb{R},$
    and the Reeb cone as
    $C_{R}:=\{\xi\in N\otimes\mathbb{R}\mid\langle\vec{m},\xi\rangle>0\quad\forall\vec{m}\in\sigma_{R}\setminus\{0\}\}.$
    We call an element $\xi$ of $C_{R}$ an abstract Reeb vector field or positive vector field. An important observation is that, for each $\xi$, we can associate a valuation ${\rm val}_{\xi}$ with the center $x\in X$ (Definition 2.10). See also there is an interpretation of $C_{R}$ as an analogue to the Kähler cone (cf., e.g., [BvC18, §2]), where it is called the Sasaki cone.
  2. ($\mathrm{ii}$)
    An affine cone 555or generalized affine cone, to clarify that it is in the broader sense than the classical case i.e., when $r=1$ and $\xi$ is regular as [Kol13b, (3.8)] is simply a triplet $(X,T\curvearrowright X,\xi)$ where $X$ is an affine algebraic $\mathbbm{k}$-scheme, $T\curvearrowright X$ is a good action, and $\xi$ is its abstract Reeb vector field. We denote the dimension of $X$ as $n$.
    $\xi$ is called regular (resp., quasi-regular) if $r=1$ and $T$ acts freely on $X\setminus\{x\}$ (resp., $r=1$). In the case when $\xi$ is regular, we denote the corresponding polarized projective scheme as $(V,L)$ so that $X=\mathop{\mathrm{Spec}}\nolimits\oplus_{m\geq 0}H^{0}(V,L^{\otimes m})$. If $\xi$ is quasi-regular, i.e., in the general $r(\xi)=1$ case, we can similarly construct the quotient $[(X\setminus x)/T]\to(X\setminus x)//T=:V$ as a projective scheme (the GIT quotient). If $X$ is normal, $V$ is also automatically normal (cf., [Mum65, Chapter I]) and we can consider the ramification index $m(D)$ for each prime divisor $D$ on $V$ and form the standard ramification divisor as a Weil $\mathbb{Q}$-divisor $\sum_{D}\frac{m(D)-1}{m(D)}D$ on $V$. This morphism $(X\setminus x)\to(V,\sum_{D}\frac{m(D)-1}{m(D)}D)$ is called the (algebraic) Seifert $\mathbb{G}_{m}$-bundle in [Kol04, Kol13b] (cf., also [LL19, §3.1]), It globally realizes the so-called transverse Kähler structure on the locally orthogonal direction to the Reeb foliation in $r=1$ case (in the irregular case, one can only locally realize it cf., [BG07]).
  3. ($\mathrm{iii}$)
    A $\mathbb{Q}$-Gorenstein (affine) cone666temporary name in this paper is (generalized) affine cone where $X$ is reduced which is normal crossing in codimension $1$, satisfying the Serre condition $S_{2}$, $K_{X}$ is $\mathbb{Q}$-Cartier, $T\curvearrowright X$ is a good action, and $\xi$ is its abstract Reeb vector field. It is called a Fano cone if $X$ is also log terminal. We also call it a (kawamata-)log-terminal cone in this paper. Similarly, if $X$ is log canonical (resp., semi-log-canonical), we call it a log canonical cone (resp., semi-log-canonical cone).
There is a valuative interpretation of positive vector field.
Definition 2.10 (Monomial valuations cf., [BFJ08], [Li18, §2.2]).
If $x\in X\curvearrowleft T$ is an irreducible affine cone, then for each positive vector field $\xi\in C_{R}$ (see Definition 2.9), we can associate a valuation ${\rm val}_{\xi}$ of $X$ centered at the vertex $x$ as follows:
(3) ${\rm val}_{\xi}(f):=\min_{\vec{m}\in\sigma_{R}}\{\langle\vec{m},\xi\rangle\mid f_{\vec{m}}\neq 0\},$
[…]
The following should be known to experts.
Lemma 2.11.
  1. ($\mathrm{i}$)
    In the setup of Definition 2.9 ($\mathrm{i}$), if $mK_{X}$ is Cartier for some positive integer $m$, then it is automatically linearly trivial.
  2. ($\mathrm{ii}$)
    For a $\mathbb{Q}$-Gorenstein affine cone $T\curvearrowright X\ni x$ with the abstract Reeb vector field $\xi$, there is a positive integer $l$ and a nowhere vanishing holomorphic section $\Omega\in\Gamma(\mathcal{O}(lK_{X}))$ and some real number $\lambda$ so that
    $L_{\xi}\Omega=\sqrt{-1}\lambda\Omega.$
    Here, $L_{\xi}$ means the Lie derivative. Moreover, $T\curvearrowright X\ni x$ with $\xi$ is a Fano cone if and only if $\lambda>0$.
Proof.
For ($\mathrm{i}$), by taking the index $1$-cover with respect to $K_{X}$, we can reduce to the quasi-Gorenstein case. By the complete reducibility of $T\curvearrowright\Gamma(mK_{X})$ we can write $K_{X}$ as a $T$-invariant divisor which does not contain the vertex, but then it is easy to see the support is empty. See [PS11, LS13] for discussions in more general case. The classical case i.e., for the regular action of $T$ with $r=1$, see [Kol13b, 3.14] for the proof of this ($\mathrm{i}$). For ($\mathrm{ii}$) is proved in [CS19, Lemma 6.1, 6.2] for the normal case. For non-normal case, the same proof works by combining with [GZ16, 16.45, 16.47 and the proof]. ∎
The last statement also follows from the Seifert $\mathbb{G}_{m}$-bundle interpretation by [Kol13b, §9.3], [LL19, §3.3.1] (cf., also [LLX18, §3.1]). $\lambda$ of the above is interestingly interpretted as log discrepancy of $X$ by C.Li [Li18]. For instance, its positivity was known to be equivalent to log terminality of $X$ as observed in [CS19, §6].
Definition 2.12.
  1. ($\mathrm{i}$)
    The multi-Hilbert function of an affine cone $T\curvearrowright X\ni x$ with the positive vector field $\xi\in N_{\mathbb{R}}$ is a map $M\to\mathbb{Z}_{\geq 0}$ defined by $\vec{m}\mapsto\operatorname{dim}R_{\vec{m}}=:\chi_{X}(\vec{m})$.
  2. ($\mathrm{ii}$)
    (cf., [MSY08, CS18, CS19]) The multi-Hilbert series of an affine cone $T\curvearrowright X\ni x$ with the positive vector field $\xi\in N_{\mathbb{R}}$  [^7] Originally called index character by [MSY08, CS18, CS19] in the context of equivariant index theory but this term may sound more familiar to algebraic geometers is defined as
    $F(\xi,t)=\sum_{\vec{m}\in M}e^{-t\langle\vec{m},\xi\rangle}\operatorname{dim}R_{\vec{m}}.$
    A priori one can regard it only as a formal function but we review in Lemma 2.14 (from [MSY06, CS18]) later that as far as $R$ is finitely generated, the series is meromorphic around 00 so that it encodes the multi-Hilbert function $\{\operatorname{dim}R_{\vec{m}}\}_{\vec{m}}$ as the Fourier coefficients. Obviously, the moment monoid, the moment cone and the Reeb cone are determined only by the multi-Hilbert function.
  3. ($\mathrm{iii}$)
    (cf., [CS18]) Take an affine cone $\tilde{T}\curvearrowright X\ni x$ with an algebraic $\mathbbm{k}$-torus $\tilde{T}(\supset T)$, its character lattice $\tilde{M}:={\rm Hom}(T,\mathbb{G}_{m})$ (resp., $\tilde{N}:={\rm Hom}(\mathbb{G}_{m},T)$)  [^8] This notation is set in this way as later we often apply to the central fiber of the test configuration (Definition 2.15 ($\mathrm{i}$)) so that $\tilde{T}$ is often $T\times\mathbb{G}_{m}$, unless it is a product test configuration. and $\eta\in\tilde{N}\otimes\mathbb{R}$. We denote the decomposition for the $\tilde{T}$-action of $R:=\Gamma(\mathcal{O}_{X})$ as $\oplus_{\tilde{m}}R_{\tilde{m}}$.
    The weighted (multi-)Hilbert series999called weighted character in [CS18] of the affine cone $\tilde{T}\curvearrowright X\ni x$ with the positive vector field $\xi\in N_{\mathbb{R}}$, for the $\eta$-direction, is defined as
    $C_{\eta}(\xi,t):=\sum_{\tilde{m}\in\tilde{M}}e^{-t\langle\vec{m},\xi\rangle}\langle\tilde{m},\eta\rangle\operatorname{dim}R_{\tilde{m}}.$
  4. ($\mathrm{iv}$)
    (cf., [HS17]) Consider any faithfully flat affine family between algebraic $\mathbbm{k}$-schemes $\pi\colon\mathcal{X}=\mathop{\mathrm{Spec}}\nolimits_{S}(\mathcal{R})\to S$ with connected $S$, with $\mathcal{O}_{S}$-algebra $\mathcal{R}$, where an algebraic $\mathbbm{k}$-torus acts on $\mathcal{X}$ fiberwise. We apply the complete reducibility of $T$ to $\mathcal{R}$ to obtain the decomposition $\mathcal{R}=\oplus_{\vec{m}\in M}\mathcal{R}_{\vec{m}}$ where $\mathcal{R}_{\vec{m}}$ denotes the $\mathcal{O}_{S}$-module corresponding to the character $\vec{m}$. If $\mathcal{R}_{\vec{m}}$ for any $\vec{m}$ is faithfully flat over $S$, we call the family $\pi\colon\mathcal{X}\to S$ is $T$-equivariantly faithfully flat over $S$, or $T$-fppf for short (if no confusion), in this paper. Note that this condition is said to be admissibility of the deformation in [HS17] and is stronger than the faithful flatness of $\pi$. See also the following Lemma 2.13 ($\mathrm{ii}$).
Here is some useful general lemma.
Lemma 2.13.
  1. ($\mathrm{i}$)
    (Characterization of good action) Take an affine algebraic $\mathbbm{k}$-scheme $X=\mathop{\mathrm{Spec}}\nolimits(R)$ on which an algebraic $\mathbbm{k}$-torus $T$ acts. It is a good action if and only if the multi-Hilbert functions are all finite i.e., $\operatorname{dim}(R_{\vec{m}})<\infty$ for any $\vec{m}\in M$, and the moment cone $\mathbb{R}_{\geq 0}\sigma_{R}$ is strictly convex i.e., does not contain any line. It is further equivalent to the non-triviality of the Reeb cone $C_{R}\neq\emptyset$.
  2. ($\mathrm{ii}$)
[…]
  3. ($\mathrm{iii}$)
[…]
Proof.
We first prove ($\mathrm{i}$). The only if direction is discussed in [LLX18, §3.1] (cf., also [LS13]). For the converse i.e., the if direction, note that $R_{\vec{0}}=\mathbbm{k}$ since it is a finite extension of $\mathbbm{k}$ and is an integral domain, while we assume $\mathbbm{k}$ is algebraically closed. Since $R_{\vec{0}}=R^{T}$ i.e., coincides with the $T$-invariant subring of $R$, it follows that the polystable locus $X^{ps}$ of $X$ consists of finite $T$-orbits. Furthermore, by the standard arguments using the Raynold operator, it follows that $X^{ps}$ is connected. Hence it consists of a single $T$-orbit, say $Tx$ for some closed point $x\in X$. Denote the identity component of the stabilizer ${\rm stab}(x)$ of $x$ as ${\rm stab}^{0}(x)$. Then the character lattice of $T/{\rm stab}^{0}(x)$ should be trivial since otherwise it would contradicts with the strict convexity of $\mathbb{R}_{\geq 0}\sigma_{R}$.
Next we prove ($\mathrm{ii}$). We can and do assume $S$ is affine and write the family as $\mathcal{X}={\rm Spec}_{R}(\mathcal{R})$ by a $(R:=)\Gamma(\mathcal{O}_{S})$-algebra. Apply the complete reducibility of $T$ to $\mathcal{R}$ to obtain $\mathcal{R}=\oplus_{\vec{m}\in M}\mathcal{R}_{\vec{m}}$, with $R$-modules $\mathcal{R}_{\vec{m}}$. Since $\mathcal{R}$ is $R$-flat, each $\mathcal{R}_{\vec{m}}$ is also $R$-flat hence locally free module because of the finitely generatedness assumption. Therefore, the assertion follows. Finally, ($\mathrm{iii}$) follows from ($\mathrm{i}$) and ($\mathrm{ii}$). ∎
We note that a certain embedded version of ($\mathrm{i}$) (resp., ($\mathrm{ii}$)) is partially proved in literature as [CS18, 4.5], [KR00, 4.1.19] (resp., [CS18, after Definition 5.1]). We also review the following for the next section.
Lemma 2.14 ([CS18, CS19]).
Recall that we set $n:=\operatorname{dim}(X)$.
  1. ($\mathrm{i}$)
    The multi-Hilbert series $F(\xi,t)$ can be written as
    $F(\xi,t)=\dfrac{A_{0}(\xi)}{t^{n}}-\dfrac{A_{1}(\xi)}{2t^{n-1}}+O(t^{2-n}),$
    with $t$ around $0\in\mathbb{C}$ where $A_{i}(\xi)(i=0,1)$ are $C^{\infty}$-function on the Reeb cone $C_{R}$ (Definition 2.9). In general, we have
    (4) $A_{0}(\xi)=\widehat{\operatorname{vol}}({\rm val}_{\xi}),$
    where the latter $\widehat{\operatorname{vol}}(-)$ means the normalized volume in the sense of [Li18] (Definition 2.27). Note that this holds for general affine cone as proved in [CS18, §4].
    Furthermore, if $r(\xi)=1$, $T$ is regular and $\xi$ is a generator of $N$, in the notation of Definition 2.9, $A_{0}(\xi)=(L^{n-1})$, $A_{1}(\xi)=(L^{n-2}.K_{V})$. Note the obvious homogeneity $A_{i}(c\xi)=c^{-n+i}A_{i}(\xi)$ for $i=0,1$.
  2. ($\mathrm{ii}$)
    The weighted multi-Hilbert series $C_{\eta}(\xi,t)$ can be written as
    $C_{\eta}(\xi,t)=\dfrac{B_{0}(\xi)}{t^{n+1}}-\dfrac{B_{1}(\xi)}{2t^{n}}+O(t^{1-n}),$
    where $B_{i}(\xi)(i=0,1)$ are $C^{\infty}$-function on the Reeb cone $C_{R}$.
    B functions are directional derivative of A functions in the sense that $B_{i}(\xi)=-D_{\eta}A_{i}(\xi)$ for $i=0,1$ where $D_{\eta}$ denotes the directional derivative along the direction $\eta$. Hence, we again have the homogeneity $B_{i}(c\xi)=c^{-n-1+i}A_{i}(\xi)$ for $i=0,1$.
Proof.
See [CS18, Theorem 3] (which depends on [KR00, §5.8]) for the proof of ($\mathrm{i}$). ($\mathrm{ii}$) follows from the derivation by terms of $\sum_{\tilde{m}\in\tilde{M}}e^{-t\langle\tilde{m},\xi-s\eta\rangle}\operatorname{dim}R_{\tilde{m}}$. See the details in the proof of [CS18, Theorem 4]. ∎
2.3. K-stability of affine cones
Definition 2.15.
  1. ($\mathrm{i}$)
    (cf., [CS18, CS19]) A (affine $T$-equivariant) test configuration of an affine cone $(T\curvearrowright X,\xi)$ means $T\times\mathbb{G}_{m}$-equivariant affine $T$-equivariantly faithfully 101010this $T$-equivariant faithful flatness is necessary e.g. to avoid $X\times(\mathbb{A}^{1}\setminus\{0\})$ and some other pathological examples, which is missed in some literature flat morphism $\pi\colon\mathcal{X}\to\mathbb{A}^{1}$ from a normal affine scheme $\mathcal{X}$, where $T$ acts only fiberwise (while acting trivially on the base $\mathbb{A}^{1}$) and $\mathbb{G}_{m}$ acts multiplicatively on the base $\mathbb{A}^{1}$ such that its restriction to $\pi^{-1}(\mathbb{A}^{1}\setminus\{0\})$ is the product $X\times(\mathbb{A}^{1}\setminus\{0\})$. A test configuration is called product test configuration if there is a $T$-equivariant isomorphism $\mathcal{X}\simeq X\times\mathbb{A}^{1}$. For simplicity, we sometimes abbreviate the whole data of a test configuration just as $\mathcal{X}$ if it does not make confusion.
  2. ($\mathrm{ii}$)
[…]
    (5) ${\rm DF}(\mathcal{X},\xi):=(n+1)A_{1}(\xi)B_{0}(\xi)-nA_{0}(\xi)B_{1}(\xi),$
    where $A_{i}(\xi)$ are that of any fibers $\pi^{-1}(s)$ for $s\in\mathbb{A}^{1}$ (see Lemma 2.13 ($\mathrm{ii}$) for the independence on $s$) $B_{i}(\xi)$ are that of $\mathcal{X}_{0}=\pi^{-1}(0)$ defined in Definition 2.12 ($\mathrm{ii}$).
    If $r=1$, $T$ is regular and $\xi$ is the generator of $N$, in the notation of Definition 2.9, we have
    (6) $A_{0}(\xi)=(L^{n-1}),$ $A_{1}(\xi)=(L^{n-2}.K_{V}),$
    (7) $B_{0}(\xi)=(\mathcal{L}^{n}),$ $B_{1}(\xi)=(\mathcal{L}^{n-1}.K_{\mathcal{V}/\mathbb{P}^{1}}),$
    where $(\mathcal{V},\mathcal{L})$ denotes the (polarized projective) test configuration in the sense of [Don02] which arise as $((\mathcal{X}\setminus\overline{\mathbb{G}_{m}\cdot(x,1)})/\mathbb{G}_{m})\to\mathbb{A}^{1}$ by [Wan12, Od13a] so that it fits to the original intersection number formula in loc.cit for the polarized projective setup.
    Then, the (generalized) affine cone $(T\curvearrowright X,\xi)$ is K-stable (resp., K-semistable [^12] Notes added: see also [Wu22, LW24] for rephrasing in more non-archimedean terminology. The latter appeared after our paper.) if and only if $\mathop{\mathrm{DF}}\nolimits(\mathcal{X},\xi)>0$ (resp., $\mathop{\mathrm{DF}}\nolimits(\mathcal{X},\xi)\geq 0$) for any non-trivial affine test configuration $\pi\colon\mathcal{X}\to\mathbb{A}^{1}$. It is said to be K-polystable if it is K-semistable and $\mathop{\mathrm{DF}}\nolimits(\mathcal{X},\xi)=0$ only occurs when $\mathcal{X}$ is a product test configuration.
  3. ($\mathrm{iii}$)
    (cf., [LX14]) A test configuration of a Fano cone is called a special test configuration if $(\mathcal{X},\mathcal{X}_{0})$ is purely log terminal.
  4. ($\mathrm{iv}$)
    (cf., [CS18, CS19]) If we allow ourselves to use the normalized volume [Li18] to be reviewed in §2.4.2, for a special test configuration $\mathcal{X}$ of a Fano cone, its Donaldson-Futaki invariant is equivalently defined as
    $\mathop{\mathrm{DF}}\nolimits(\mathcal{X},\xi)=\frac{d}{dt}|_{t=0}\widehat{\operatorname{vol}}_{\mathcal{X}_{0}}({\rm val}_{\xi-t\eta}),$
    up to a dimensional constant $2((n+1)!(n-1)!)$. where $\eta$ is the holomorphic vector field induced by the $\mathbb{G}_{m}$-action on the central fiber $\mathcal{X}_{0}$ of $\mathcal{X}$. Note also that if we normalize $\xi$ and $\eta$ i.e., to multiply suitable positive real number constants, so that $\xi-\epsilon\eta$ for $\epsilon\ll 1$ all satisfy the gauge fixing condition ([CS19, Definition 6.3, Proposition 6.4]) i.e., replace $\eta$ by $T_{\xi}(\eta):=\frac{A(\xi)\eta-A(\eta)\xi}{n}$ (cf., [LLX18, 3.9]), we can replace the above $\mathop{\widehat{\mathrm{vol}}}\nolimits$ by the (unnormalized) volume function $\operatorname{vol}(-)$ i.e., consider instead
    $D_{-T_{\xi}(\eta)}\operatorname{vol}_{\mathcal{X}_{0}}(\xi)$
    as it has the same sign as ${\rm DF}(\mathcal{X})$ where $D_{-T_{\xi}}(-)$ again denotes the directional derivative. One of its benefits is that we know the strict convexity of $\operatorname{vol}(-)$ on $C_{R}$ by [MSY08] (later generalized by [LX18, §3.2.2] to possibly singular T-varieties using the Okounkov body [Oko96, LM09]).
Example 2.16 (From one parameter subgroup to test configuration).
If we consider a multi-Hilbert scheme $H$ for affine closed subscheme of $\mathbb{A}^{N}$ of positive weights for an algebraic $\mathbbm{k}$-torus $T$ with fixed multi-Hilbert function, the centralizer of $T$ in ${\rm GL}(N)$ which we denote as $G$, naturally acts on $H$. If we take one parameter subgroup $\rho\colon\mathbb{G}_{m}\to G$ and a point $[X]\in H$, the family over $\overline{\rho(\mathbb{G}_{m})\cdot[X]}$ is automatically a test configuration. This is essentially the way [CS18, Definition 5.1] first defined the test configurations. Indeed, it is also easy to see that all test configuration, in our above more abstract sense, arises in this manner.
Note that more Donaldson-type (i.e., [Don02]) definition is also proved to be equivalent, as an analogue of [LX14], by Wu [Wu23]. There is also a partial related result in [LWX21, 4.3], which we extend in the proof of Theorem 3.6.
We also prepare the Duistermaat-Heckman measures and norm functionals analogous to the global projective setup, following [DH82, Od12a, His16, BHJ17, Der16, Wu22].
Definition 2.17.
  1. ($\mathrm{i}$)
    ([DH82, His16]) For a $\mathbb{Q}$-Gorenstein affine cone $T\curvearrowright X\ni x$ with an additional commuting action of $\mathbb{G}_{m}$, we define the Duistermaat-Heckman measure $DH(X)$ for the $\mathbb{G}_{m}$-action as the probability measure
    $\lim_{c\to\infty}\biggl{(}\sum_{\begin{subarray}{c}\lambda\in\mathbb{Z},\\ \vec{m}\in M,\langle\vec{m},\xi\rangle<c\end{subarray}}\dfrac{\operatorname{dim}(R_{\vec{m}})_{\lambda}}{\operatorname{dim}(R_{\vec{m}})}\quad\delta_{\frac{\lambda}{\langle m,\xi\rangle}}\biggr{)}.$
    Here, $(R_{\vec{m}})_{\lambda}$ is the $\mathbbm{k}$-linear subspace of $R_{\vec{m}}$ with the $\mathbb{G}_{m}$-weight $\lambda$ and $\delta_{a}$ of $a\in\mathbb{R}$ denotes the Dirac measure supported on $a\in\mathbb{R}$. Note that the existence of the limit measure (convergence) easily follows from approximating $\xi$ by rational vectors, to which [His16, BHJ17] applies.
    The Duistermaat-Heckman measure of a ($T$- equivariant affine) test configuration $\pi\colon\mathcal{X}\to\mathbb{A}^{1}$ refers to that of $\mathbb{G}_{m}\curvearrowright\mathcal{X}_{0}=\pi^{-1}(0)$.
  2. ($\mathrm{ii}$)
[…]
Lemma 2.18 (cf., [Od12a, 2.7(i), p.2283 l2], [BHJ17, §7.2] [Der16, 3.10, 3.11, §4]).
The strict positivity $(I^{\rm NA}-J^{\rm NA})(\mathcal{X})>0$ holds unless $\mathcal{X}$ is the trivial test configuration i.e., $(T\times\mathbb{G}_{m})$-equivariantly isomorphic to $X\times\mathbb{A}^{1}$.
Proof.
By approximating $\xi$ by a sequence of rational elements in $N_{\mathbb{Q}}$, due to [BHJ17, 7.8], it is enough to show that $J^{\rm NA})(\mathcal{X})>0$ unless $\mathcal{X}$ is the trivial test configuration. Furthermore, again by the same approximation, [BHJ17, 7.8] implies that $J^{\rm NA}(\varphi_{\mathcal{X}})=\sup{\rm supp}DH(\mathcal{X})-\int_{\mathbb{R}}\lambda DH_{\mathcal{X}}$. If this attains 00, it follows that $DH_{\mathcal{X}}$ is a Dirac measure whose support is a single point. Since we assume $\mathcal{X}$ is normal, it is then trivial test configuration. See also [Der16, 4.7]. ∎
Problem 2.19 (cf., [Od13b]).
Can we prove that any K-semistable $\mathbb{Q}$-Gorenstein affine cone $(T\curvearrowright X,\xi)$ is semi-log-canonical by extending the method to [Od13b]? What is the differential geometric meaning behind it if it is true? (cf., e.g. [BG14]).
The following is the Yau-Tian-Donaldson type correspondence, which generalizes the case of Fano manifolds mainly after [Ber16, DaSz16, CDS15, Tia15].
Theorem 2.20.
  1. ($\mathrm{i}$)
    ([CS18, CS19]) A (smooth) Fano cone $T\curvearrowright X$ admits a structure of Ricci-flat Kähler metric, which is a cone metric of some Sasaki metric on the link, with metric Reeb vector field $\xi$ if and only if $(T\curvearrowright X\ni x,\xi)$ is K-polystable in the sense of Definition 2.15.
  2. ($\mathrm{ii}$)
    ([Li21]) A (possibly klt) Fano cone $T\curvearrowright X$ admits a structure of Ricci-flat (weak) Kähler metric, which is a cone metric of some Sasaki metric on the smooth locus of the link, with metric Reeb vector field $\xi$ if and only if $(T\curvearrowright X\ni x,\xi)$ is K-polystable in the sense of Definition 2.15.
Such a cone type complete Ricci-flat Kähler metric $g$ on $(T\curvearrowright X\ni x,J,\xi)$ is often called Ricci-flat Kähler cone metric and such a Fano cone with Ricci-flat Kähler cone metric (cf., Prop 2.3) is often called Calabi-Yau cone in the literatures under some additional smoothness assumptions, which is in particular Fano cones. We call $\mathop{\widehat{\mathrm{vol}}}\nolimits(x\in X)$ its volume density (cf., Proposition 2.7 for the motivation).
We need to be careful not to mix up with log canonical cone we define in Definition 2.9, which are often the affine cone over polarized projective (log canonical) Calabi-Yau varieties (see [Kol13b, 3.1]). We also note the latter singular version ($\mathrm{ii}$) relies on the recent theory of “weighted” framework by using the moment maps (cf., [Ino19a, Lah19, Ino19b, HaL23, AJL23, Li21]).
Motivated by the above Definition 2.20 ($\mathrm{ii}$), we introduce the following terminology, allowing singularities to extend the notions reviewed in §2.1.
Definition 2.21.
A (compact) Sasaki-Einstein space (or Sasaki-Einstein manifold with singularities) is a metric space which is decomposed as $S=S^{\rm reg}\sqcup S^{\rm sing}$ (an open dense subset which is a manifold and closed singular locus) with an Einstein (Riemannian) metric $g$ on $S^{\rm reg}$ which induces the metric structure, together with a Reeb vector field $\xi\in\Gamma(T_{S^{\rm reg}})$, such that the following holds:
corresponding complex structure extends to the cone $C(S):=(S\times\mathbb{R}_{>0})\cup 0$ which underlies a Ricci-flat weak Kähler cone (in the sense of [Li18] and the previous Theorem 2.20 ($\mathrm{ii}$)).
We hope for more intrinsic equivalent definition.
2.3.1. CM $\mathbb{R}$-line bundle
The CM line bundle, first introduced by [FS90, §10, 10.3-5] and generalized by [PT06, FR06] to singular setup, gives a canonical ample line bundle on the K-moduli of canonically polarized varieties ([Fjn18, KP17, PX17]), polarized Calabi-Yau varieties ([Vie95, Od21]), Fano varieties ([CP21, XZ20]) as we expect the same for more general polarized K-stable polarized varieties (cf., [FS90, Od10]).
As in the setup for flat families of the polarized varieties, we first introduce CM $\mathbb{R}$-line bundle on the moduli stack and then discuss its descended $\mathbb{R}$-line bundle on the good moduli space. Note that it is originally motivated by the generalization theory of Weil-Petersson metrics [FS90], which was originally obtained as a Quillen-type metric for virtual vector bundles à la Donaldson.
Definition 2.22.
Consider an arbitrary flat family of $n$-dimensional Fano cones $T\curvearrowright\mathcal{X}\to S$, whose relative coordinate ring we decompose as $\mathcal{X}={\rm Spec}_{S}\oplus_{\vec{m}}\mathcal{R}_{\vec{m}}$. Here, $\vec{m}$ runs over the character lattice $M:={\rm Hom}(T,\mathbb{G}_{m})$. Then, we consider the following series
(8) $L(\xi,t):=\bigotimes_{\vec{m}\in M}(\operatorname{det}\mathcal{R}_{\vec{m}})^{\otimes e^{-t\langle\vec{m},\xi\rangle}},$
which give $\mathbb{R}$-line bundle i.e., element of ${\rm Pic}(S)\otimes_{\mathbb{Z}}\mathbb{R}$ for each complex number $t$ with ${\rm Re}(t)>0$ (the same proof as [CS18, 4.2]). Then, by the same arguments as [CS18, 4.3], it has following expansion of meromorphic type:
(9) $L(\xi,t)=\frac{\mathcal{B}_{0}n!}{t^{n+1}}+\frac{\mathcal{B}_{1}(n-1)!}{t^{n}}+O(t^{1-n}),$
if we write in the additive notation (note that ${\rm Pic}(S)\otimes\mathbb{R}$ is a finite dimensional real vector space). After the following normalization
(10) $\mathcal{C}_{0}(\xi)$ $:=n!\mathcal{B}_{0},$
(11) $\mathcal{C}_{1}(\xi)$ $:=-\frac{(n-1)\cdot(n!)}{2}\mathcal{B}_{0}-(n-1)!\mathcal{B}_{1},$
we define the CM $\mathbb{R}$-line bundle $\lambda_{\rm CM}(T\curvearrowright\mathcal{X})$ as $-(n-1)A_{1}(\xi)\mathcal{C}_{0}+nA_{0}(\xi)\mathcal{C}_{1}$ as an element of ${\rm Pic}(S)\otimes\mathbb{R}$. Note that the above normalization is suitable in the sense that if $\xi$ is regular and $S$ is a smooth proper curve,
(12) $\operatorname{deg}\mathcal{C}_{0}(\xi)$ $=(\mathcal{L}^{n}),$
(13) $\operatorname{deg}\mathcal{C}_{1}(\xi)$ $=(\mathcal{L}^{n-1}.K_{\mathcal{Y}/S}),$
where $\mathcal{Y}\to S$ denotes the $\xi$ quotient of $\mathcal{X}\setminus\sigma(S)$ where $\sigma$ denote the vertex section, and $\mathcal{L}$ is the corresponding ample line bundle.
See also a related general construction in [Ino20].
2.4. Donaldson-Sun theory and normalized volume
2.4.1. Original work of Donaldson-Sun
In [DS17], for log terminal singularities with (singular) Kähler-Einstein metrics, the analytic and even algebraic nature of the metric tangent cone is started to explore. This is somewhat a local analogue of [DS14]. As a technical assumption, [DS17] assumes the metrized log terminal singularities appear in the non-collapsed Gromov-Hausdorff limits (polarized limit space) of Kähler-Einstein manifolds. After [DS17], more algebraic works by [Li18, LL19, CS18, CS19] appear, which refines and more algebraize the conjectural picture which is now a theorem by [LX20, LX18, LWX21, BLQ22, XZ22, XZ21] (cf., the surveys [LLX18, Zhu23b]).
We briefly review this whole story in this subsection. In the work of [DS14, DS17], the following setup is mainly considered (although being somewhat more general).
Setup 1.
Consider a flat projective family of $n$-dimensional smooth polarized varieties $f\colon(\mathcal{X},\mathcal{L})\to B$ over a connected base algebraic $\mathbbm{k}$-scheme $B$ with a section $\sigma\colon B\to\mathcal{X}$ which satisfies $K_{\mathcal{X}}\equiv_{B}\mathcal{L}^{\otimes k}$ with a constant $k\in\mathbb{Q}$ 131313[DS14, DS17] only requires to fix $n$ and the constancy of the volume of the fiber $\mathcal{L}_{b}^{n}$ for each $b\in B$. Take a sequence of points $b_{i}\in B(i=1,2,\cdots)$ whose $f$-fibers are $(p_{i}=\sigma(b_{i})\in X_{i},L_{i})$ and admit Kähler-Einstein metrics with the Kähler class in $2\pi c_{1}(L_{i})$.
We take a sequence of compact Kähler-Einstein manifolds $(p_{i}=\sigma(b_{i})\in X_{i},L_{i},g_{i},\omega_{i})_{i=1,2,\cdots}$ which satisfies the non-collapsing condition:
there is a positive constant $c$ such that if we consider the geodesic ball $B_{r}(p_{i})$ in $X_{i}$ with center $p_{i}=\sigma(b_{i})$ of radius $r\in[0,1]$ we have 141414(or equivalently the same for $r$ in $[0,\epsilon]$ for small fixed $\epsilon$ independent of $i$)
(14) $\operatorname{vol}(B_{r}(p_{i}))\geq cr^{2n}\text{ for }i\gg 0.$
Following may be worth noted.
Lemma 2.23 (non-collapsing condition as log-terminality).
If $f$ extends to a locally stable projective family $\overline{f}\colon\overline{\mathcal{X}}\to\overline{B}$ whose base $\overline{B}$ contains $B$ as a Zariski open subset and restricts to $f$ over $B$ i.e., $f^{-1}(B)=\mathcal{X}$, and $b_{i}$ converges to a point $b\in\overline{B}$. We further assume that $f^{-1}(b)$ is K-polystable. Then, the above non-collapsing condition (14) holds if and only if $\sigma(b)\in f^{-1}(b)$ is log terminal.
Proof.
The only if direction follows from [DS14] and the characterization of log terminality as local volume boundedness of the adapted measure in [EGZ09]. If $k<0$, where $k$ is as in the Setup 1 above, then both conditions automatically hold. In the case $k=0$, the if direction also follows from [EGZ09, DGG23] (see also [Od21]). In the case $k>0$, it follows from [BG14, Son17, DGG23]. ∎
Theorem 2.24 ([DS17]).
  1. ($\mathrm{i}$)
    (loc.cit 1.1) Passing to a subsequence of $i$, there is a polarized limit space151515the enhanced notion of pointed Gromov-Hausdorff limit to consider complex structure as defined in [DS14, DS17] $Z\ni p$ with a singular Kähler metric $g$ as a complex analytic space.
  2. ($\mathrm{ii}$)
    (loc.cit 1.3) Suppose we take a sequence of $(p\in Z,ag)$ with $a\to\infty$, then there is a unique polarized limit space as a log terminal affine variety $C$ with Ricci-flat Kähler cone (singular) metric on which a positive dimensional algebraic $\mathbb{C}$-torus $T(\mathbb{C})$-acts.
  3. ($\mathrm{iii}$)
    (loc.cit §2.3, §3) The above construction of $C$ from $Z\ni p$ can be separated into $2$-steps i.e., $(Z\ni p)\rightsquigarrow W\rightsquigarrow C$, where $W$ is a Fano cone.
The above $C$ is called (local) metric tangent cone of $p\in Z$ with respect to $g$. There are also related results in [VdC11]. The above item ($\mathrm{ii}$) above inspired the later developments in particular by the following conjecture.
Conjecture 2.25 ([DS17, 3.22]).
For any log terminal analytic space $x\in X$ with a weak Kähler-Einstein metric, the metric tangent cone only depends on the complex analytic germ of $x\in X$ (or $\widehat{\mathcal{O}_{X,x}}$, equivalently).
In the next subsection, we also review the developments after the above conjecture.
2.4.2. Normalized volume [Li18]
Motivated by the above Donaldson-Sun theory, there were developments of algebro-geometric machinery which in particular partially solved the above conjecture under Setup 1. Here we review the developments. Henceforth, we fix a closed point $x\in X$ as a $n$-dimensional klt singularities over $k$, on which we put the trivial valuation. We consider the space ${\rm Val}_{k}(\widehat{\mathcal{O}_{x,X}})$ of valuations $v$ of the completed stalk $\widehat{\mathcal{O}_{X,x}}$ whose center is the maximal ideal $\mathfrak{m}_{X,x}$. Note that it in particular induces a valuation of $X$ centered at $x$. The space of such valuations are denoted by ${\rm Val}_{x}(X)$.
Definition 2.26 (Local volume).
([ELS03], [LM09]) For each $v\in{\rm Val}_{k}(\widehat{\mathcal{O}_{x,X}})$, we define the following.
${\rm vol}_{x}(v):=\limsup_{m\to\infty}\dfrac{\operatorname{dim}_{k}(\mathcal{O}_{X,x}/\{f\mid v(f)\geq m\})}{m^{n}/n!}$
[…]
In the case $X$ has a structure of affine cone and $v$ comes from a positive vector field, recall that Lemma 2.14 (4) (from [MSY06, CS18]) gives another expression of this (unnormalized) volume function $\operatorname{vol}(-)$ in terms of multi-Hilbert function.
On the other hand, we take the subspace of the space ${\rm Val}_{x}(X)$ (resp., ${\rm Val}_{k}(\widehat{\mathcal{O}_{x,X}})$) defined by $A_{X}(v)=n$ (cf., [MSY08, CS19]) and write as ${\rm Val}_{x}^{n}(X)$ (resp., ${\rm Val}_{k}^{n}(\widehat{\mathcal{O}_{x,X}})$).
Definition 2.27 ([Li18]).
Define the normalized volume of $x\in X$ for $v\in{\rm Val}_{k}(\widehat{\mathcal{O}_{x,X}})$ as
(15) $\widehat{{\rm vol}}_{x}(v):=A_{X}(v)^{n}{\rm vol}_{x}(v).$
Here, if $A_{X}(v)=+\infty$, then we also set $\widehat{{\rm vol}}_{x}(v):=+\infty$. Equivalently, we consider ${\rm vol}_{x}(v)$ on the normalized section ${\rm Val}_{k}^{n}(\widehat{\mathcal{O}_{x,X}})$.
By the effect of normalization using the log discrepancy $A_{X}$, for any positive real number $\lambda$,
(16) $\widehat{\operatorname{vol}}(\lambda v)=\widehat{\operatorname{vol}}(v)$
so that the normalized volume gives a function
(17) $\widehat{{\rm vol}}\colon({\rm Val}_{k}(\widehat{\mathcal{O}_{x,X}})\setminus\{v_{\rm triv}\})/\mathbb{R}_{>0}\to\mathbb{R}\cup\{+\infty\}.$
Here, $v_{\rm triv}$ is the trivial valuation and the action of $\mathbb{R}_{>0}$ on $({\rm Val}_{k}(\widehat{\mathcal{O}_{x,X}})\setminus\{v_{\rm triv}\})$ is simply given by the multiplication on the real valuations. The quotient
(18) $({\rm Val}_{k}(\widehat{\mathcal{O}_{x,X}})\setminus\{v_{\rm triv}\})/\mathbb{R}_{>0}$
or homeomorphic ${\rm Val}_{k}^{n}(\widehat{\mathcal{O}_{x,X}})$ is often called non-archimedean link after the topological theory of links of complex algebraic singularities ([Sas60, Bri66, Mil69, BG07]). We note that clearly the latter normalization convention is motivated by the Theorem 2.8 ($\mathrm{ii}$).
Remark 2.28.
We also note that if $v$ comes from the plt blow up $\pi\colon X^{\prime}\to X$ with the exceptional divisor (Kollár component161616in the sense of [LX18] henceforth, see also the original [Sho96, 3.1], [Pro00, 2.1]) $E$, $\mathop{\widehat{\mathrm{vol}}}\nolimits(v)$ equals the generalized log Donaldson-Futaki invariant ${\rm DF}(X^{\prime}\to X,-(K_{X^{\prime}/X}+E))$ defined in [Od15b]. In the theory, blow up $\pi$ is regarded as an analogue of test configuration in the sense of [Don02].
2.5. Donaldson-Sun $2$-step degeneration and its algebraization
2.5.1. General procedure
By making use of the theory of normalized volumes in the previous section, the Donaldson-Sun $2$-step degeneration theory is now put in a format of algebraic geometry and developped, due to the work of [Li18, LL19, Blu18, CS18, CS19, LX18, LWX21, Xu20, XZ22, XZ21, BLQ22]. We briefly summarized the conclusion as follows.
Theorem 2.29.
Suppose $\mathbbm{k}$ is algebraically closed field of characteristic 00. We fix a pointed $n$-dimensional log-terminal $\mathbbm{k}$-variety $x\in X$, consider the stalk $\mathcal{O}_{X,x}$ and completed stalk $\widehat{\mathcal{O}_{X,x}}$ (Only the formal germ of $x$ matters below, hence $X$ is not assumed to be proper in general). Then we have the following.
  1. ($\mathrm{i}$)
    ([Blu18, Xu20, XZ21, BLQ22]) There is a unique (necessarily quasi-monomial [Xu20]) valuation $v=v_{X}\in{\rm Val}_{k}(\widehat{\mathcal{O}_{x,X}})\subset{\rm Val}_{x}(X)$ up to the $\mathbb{R}_{>0}$-action, which minimizes 171717analogue and the generalization of the volume minimization in [MSY06, MSY08] (called “Z-minimization” in loc.cit) the normalized volume $\mathop{\widehat{\mathrm{vol}}}\nolimits(-)$. We denote the rank of (the groupification $M$ of) $\Gamma:={\rm Im}(v_{X})$ as $r$ and the dual lattice as $N:={\rm Hom}(M,\mathbb{Z})$.
  2. ($\mathrm{ii}$)
    ([LX20, LX18, XZ22]) ${\rm gr}_{v}(\mathcal{O}_{X,x}):=\oplus_{s\in\Gamma}(\{f\in\mathcal{O}_{X,x}\mid v(f)\geq s\}/\{f\in\mathcal{O}_{X,x}\mid v(f)>s\})$ is of finite type over $\mathbbm{k}$ ([XZ22] for general $r\geq 1$ case). There is a natural action of the algebraic $\mathbbm{k}$-torus $T:=N\otimes\mathbb{G}_{m}$ on $W:=\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v}(\mathcal{O}_{X,x})$ and this provides K-semistable Fano cone structure to $W$ ([LX20, LX18]). We denote its vertex as $x_{W}$.
  3. ($\mathrm{iii}$)
    ([LWX21]) There is a (non-canonical) affine test configuration of $W$ to a unique K-polystable Fano cone $C$. In particular, $C$ is uniquely determined by germ of $x\in X$ (i.e., Conjecture 2.25 by Donaldson-Sun [DS17] holds in the case of Setup 1).
  4. ($\mathrm{iv}$)
    ([DS14, CS18, LWX21, Li21]) If $\mathbbm{k}=\mathbb{C}$, $C$ has a (weak) Ricci-flat Kähler cone metric. Further, if $X$ is compactified to the (non-collapsed) polarized limit space of Kähler-Einstein manifolds in the sense of [DS14], [DS17, §2.1], then $C$ is a metric tangent cone of $X\ni x$ and is unique.
2.5.2. Local vs global volume
The following type of inequalities are very important to study K-stability of singular $\mathbb{Q}$-Fano varieties. In the dimension $2$, [OSS16] originally used the Bishop-Gromov type inequality to deduce a similar (weaker) result which we also review.
Theorem 2.30 ([Liu18]).
For any K-semistable $\mathbb{Q}$-Fano variety $X$ and $x$ its closed point, we have
(19) $(-K_{X})^{n}\leq\biggl{(}1+\frac{1}{n}\biggr{)}^{n}\widehat{\rm vol}(v).$
Here is a weaker version, which in turn is a straightforward consequence of the Bishop-Gromov inequality for orbifolds.
Theorem 2.31 (cf., [OSS16]).
If there is a quotient singularity of the type $\mathbb{C}^{n}/\Gamma$ ($\Gamma$ is a finite subgroup of ${\rm GL}(n,\mathbb{C})$) on a Kähler-Einstein $\mathbb{Q}$-Fano variety $X$, we have
(20) $(-K_{X})^{n}\leq\frac{1}{|\Gamma|}\cdot\dfrac{2(2n-1)^{n}\cdot n!}{(2n-1)!!=(2n-1)\cdot(2n-3)\cdots 3\cdot 1}.$
The estimation comes from the comparison of the (real) space form, rather than the complex space form $\mathbb{P}^{n}$, on which we can not expect (complex) algebraic variety structure in general.
Proof.
The proof in the $2$-dimensional case ([OSS16, 2.7] cf., also [Tia90]181818cf., Remark 2.8 of [OSS16]) naturally extends. Apply (orbifold version of) the Bishop-Gromov comparision theorem to $X$ to obtain
$\frac{(2\pi)^{n}(-K_{X})^{n}/n!}{\operatorname{vol}(S^{2n}(2n-1))}\leq\frac{1}{|\Gamma|}.$
Hence the assertion follows from
$\operatorname{vol}(S^{2n}(2n-1))$ $=(2n-1)^{n}\cdot\frac{(2n+1)\pi^{n+\frac{1}{2}}}{\Gamma(n+\frac{3}{2})=(n+\frac{1}{2})(n-\frac{1}{2})\cdots\frac{1}{2}\times\sqrt{\pi}}$
$=(2n-1)^{n}\cdot\frac{2\pi^{n}}{(n-\frac{1}{2})(n-\frac{3}{2})\cdots\frac{1}{2}}.$
2.5.3. Toric case
We can see the strength of the K-stability theory and Donaldson-Sun theory by one of simplest example - toric case. As first pointed out by [CS19, §1], it readily re-proves a somewhat weaker version of the following celebrated result by Futaki-Ono-Wang [FOW09] and its generalization to singular case [Ber20]. Note that the condition in loc.cit Theorem 1.2, the meaning of the assumptions in loc.cit Theorem 1.2 (on the vanishing of the de Rham cohomology class of the contact bundle and the positivity basic first Chern class of the normal bundle of the Reeb foliation) is later clarified to be equivalent to the ($\mathbb{Q}$-)K-triviality which matches Definition 2.9 and Lemma 2.11. Hence, we restate their result in that manner, although in a somewhat weak statements (on the stability side).
Theorem 2.32 (cf., [FOW09], [Ber20]).
Suppose $T\curvearrowright X=C(S)$ is a toric Fano cone i.e., a (log terminal) Fano cone which admits an action of algebraic $\mathbbm{k}$-torus $T^{\prime}$ which preserves the structure. We denote $N^{\prime}:={\rm Hom}(\mathbb{G}_{m},T^{\prime})$ and $M^{\prime}:={\rm Hom}(T^{\prime},\mathbb{G}_{m})$. Then $X$ is K-semistable and $T^{\prime}$-equivariantly K-polystable as Fano cone with respect to a (unique) Reeb vector field $\xi\in N^{\prime}\otimes\mathbb{R}$.
Proof.
The (unique) $\mathop{\widehat{\mathrm{vol}}}\nolimits$-minimizing valuation $v=v_{X}$ of $\mathcal{O}_{X,x}$ is $T^{\prime}$-invariant hence is a toric valuation in the sense of [Blu18, §8]. Thus, if we write $C(S)=\mathop{\mathrm{Spec}}\nolimits R=\oplus_{\vec{m}\in M^{\prime}}R_{\vec{m}}$ for $M^{\prime}:={\rm Hom}(T^{\prime},\mathbb{G}_{m})$ as the decomposition of the coordinate ring by the $T$-action, it is easy to see ${\rm gr}_{v}(R)=R$ so that $W=X$ in the Theorem 2.29. In particular, $X$ is K-semistable with respect to a Reeb vector field in $N\otimes\mathbb{R}$ and $T\subset T^{\prime}$. Further, if there is a $T^{\prime}$-equivariant special test configuration $(\mathcal{X},\xi,\eta)$, then since $\operatorname{dim}(X)=\operatorname{dim}(T^{\prime})$, $T\times\mathbb{G}_{m}$ can not act effectively on any component of $\mathcal{X}_{0}$ so that $\mathcal{X}$ is a $T^{\prime}$-equivariant product test configuration. Hence, because $\mathcal{X}_{0}$ has vanishing Futaki invariant with respect to $\xi$ by its K-semistability, it follows that $X$ is $T^{\prime}$-equivariantly K-polystable with respect to $\xi^{\prime}$. ∎
This contrasts with the fact that toric Fano manifolds do not necessarily have Kähler-Einstein metrics, but this is due to the flexibility of the (abstract) Reeb vector fields.
2.6. Affine generalized test configurations vs filtered blow ups
As a preparatory general material, we excerpt a part of [Od24b] on the theory of generalized test configurations and give an alternative description of affine generalized test configuration for general affine varieties (cf., also [Ino22, BJ21]). In short, generalized test configurations correspond to certain ideals sequences or their (filtered) blow ups. This generalizes the perspective of [Od13a].
Firstly, we recall the definition of generalized test configurations.
Definition 2.33.
For an affine toric variety $U_{\tau}$ which corresponds to a rational polyhedral cone $\tau\subset N\otimes\mathbb{R}=:N_{\mathbb{R}}$ as in Notation 1, a generalized test configuration of affine variety $X$ is a faithfully flat $T$-equivariant affine morphism $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$ with general fiber $X$.
If general fibers are affine cones in the sense of Definition 2.15 with respect to a $p$-fiberwise action of an additional algebraic $\mathbbm{k}$-torus $T^{\prime}$, we further require the following:
$\mathcal{Y}=\mathop{\mathrm{Spec}}\nolimits\mathcal{R}$ with a finite type $\Gamma(U_{\tau})$-algebra $\mathcal{R}$ that decomposes as $\mathcal{R}=\oplus_{\vec{m}\in M}\mathcal{R}_{\vec{m}}$, to the $T^{\prime}$-eigen-$\Gamma(U_{\tau})$-submodules, then $\mathcal{R}_{\vec{m}}$ for each $\vec{m}$ is a locally free module.
Example 2.34.
For a klt affine variety $X\ni x$, take the valuation $v_{X}$ of minimizing normalized volume ([Li18, Blu18]) and set $W$ as $\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v_{X}}(\mathcal{O}_{X,x}))$. Denote the groupificiation $N$ of ${\rm Im}(v_{X})$ and a rational polyhedral cone $\tau\subset N_{\mathbb{R}}$ which includes $\xi$.
Naturally, there is a canonical generalized test configuration $\pi_{\tau}\colon\mathcal{X}_{\tau}\twoheadrightarrow U_{\tau}$ of $X$ whose fiber over $p_{\tau}$ is $W$ ([Tei03, LX18]). This gives a positive weight deformation of $W$ in the sense of [Od24b]. See more details in [Od24b, §2] (cf., also [LX18]).
Lemma 2.35.
Fix an affine $\mathbbm{k}$-variety $\mathop{\mathrm{Spec}}\nolimits(R)=X$ and a rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$ as before. Then, the following two sets have natural bijective correspondence with each other:
  1. ($\mathrm{i}$)
    the set of isomorphic classes of (affine) generalized test configuration over $U_{\tau}$ which dominates 191919analogous to the blow up type (semi) test configurations in [Od13b] $X\times U_{\tau}$, denoted by $\pi\colon\mathcal{X}=\mathop{\mathrm{Spec}}\nolimits(\mathcal{R})\twoheadrightarrow\mathop{\mathrm{Spec}}\nolimits(\mathbbm{k}[\mathcal{S}_{\tau}])$ (here, $\mathcal{S}_{\tau}:=\{x\in M\otimes\mathbb{R}\mid\langle x,\tau\rangle\subset\mathbb{R}_{\geq 0}\}\cap M$)
  2. ($\mathrm{ii}$)
    the set of sequence of ideals of $R$ which we write $\{I_{\vec{m}}\subset R\}_{\vec{m}\in M}$ (or $\vec{m}\mapsto I_{\vec{m}}$) that satisfies the following:
    • $I_{\vec{m}}\cdot I_{\vec{m}^{\prime}}\subset I_{\vec{m}+\vec{m}^{\prime}}$ for any $\vec{m},\vec{m}^{\prime}\in M$,
    • $I_{\vec{m}}\subset I_{\vec{m}^{\prime}}$ for any $\vec{m},\vec{m}^{\prime}\in M$ with $\vec{m}^{\prime}-\vec{m}\in\mathcal{S}_{\tau}$,
    • and $\oplus_{\vec{m}}I_{\vec{m}}\cdot\chi(\vec{m})(=\mathcal{R})$ is of finite type over $\mathbbm{k}$ (or equivalently, over $\mathbbm{k}[\mathcal{S}_{\tau}]$),
    • $I_{\vec{m}}\neq R$ for (some or any) $\vec{m}\in-\mathcal{S}_{\tau}^{o}$.
[…]
(21) $\mathfrak{a}_{\xi,k}:=\sum_{\vec{m},\langle\vec{m},\xi\rangle\leq-k}I_{\vec{m}}.$
[…]
  1. (3)
    the relative isomorphism class (over $X$) of a birational projective morphism $h\colon Y_{\xi}\to\mathop{\mathrm{Spec}}\nolimits(R)$ with an effective $h$-antinef $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor $E_{\xi}$.
Further, if we consider the base change of $\pi$ via the toric morphism $\mathbb{A}^{1}\to U_{\tau}$ which corresponds to $l\xi$ and denote as $\mathcal{X}_{l\xi}\to\mathbb{A}^{1}$, $(\mathcal{X}_{l\xi},(\mathcal{X}_{l\xi})_{0})$ is log canonical if and only if $(Y_{\xi},E_{\xi})$ is log canonical.
Proof.
We first prove the bijection between the first two i.e., ($\mathrm{i}$) and ($\mathrm{ii}$). Since $\mathcal{R}$ is flat over $\mathbbm{k}[\mathcal{S}_{\tau}]$, $\mathcal{R}\subset\mathcal{R}\otimes_{\mathbbm{k}[\mathcal{S}_{\tau}]}\mathbbm{k}[M]\simeq R[M]$. We put $I_{\vec{m}}:=\{f\in R\mid\chi(\vec{m})f\in\mathcal{R}\}$ for each $\vec{m}\in M$. The condition that $\mathcal{X}$ dominates $X\times U_{\tau}$ ensures that $I_{\vec{m}}$ are ideals of $R$ and the last condition of $I_{\vec{m}}$ in the item ($\mathrm{ii}$) is equivalent to the surjectivity of $\pi$ i.e., $\mathcal{R}\neq(\oplus_{\vec{m}\in\mathcal{S}_{\tau}\setminus 0}R)\mathcal{R}$.
Next we discuss the map from ($\mathrm{ii}$) with normal $\mathcal{R}$ to the set (3). Fix $\xi$ and consider its multiple $l\xi\in N$. The proof of [LWX21, 2.20, 2.21] applies to obtain a blow up $h\colon Y_{\xi}\to X$. Further this construction naturally associates an effective $h$-antiample $\mathbb{Q}$-divisor $E_{\xi^{\prime}_{q}}$. By loc.cit, the last assertion of the equivalence of log canonicities also follows. We complete the proof. ∎
2.7. (Higher) $\Theta$-stratification
One of the main tools in the algebraic construction of K-moduli of Fano varieties (cf., [Xu20, BHLLX21, LXZ22]) is stack-theoretic incarnation and generalization of the theory of Harder-Narasimhan filtration, which is introduced by [HL14] and called the $\Theta$-stratification. Roughly put, it allows a family-wise optimal destabilization along one parameter family. Correspondingly, the properness-type criterion after Langton [Lan75] is established in [AHLH23]. Because of the irrational nature of $\mathop{\widehat{\mathrm{vol}}}\nolimits$-minimizing valuations (associated to the Reeb vector fields), it is more convenient and suitable to use a generalization of [HL14, AHLH23] to multi-variable parameter family setup and degenerations in its irrational direction, which is done as a part of a theory of [Od24b, §2, Ex 2.10]. We briefly review some of the main definitions and theorems below, in a simplified weaker form, following the toric Notations 1.
Definition 2.36 ([HL14, Od24b]).
Generalizing [HL14] ($r=1$ case), a higher $\Theta$-stratum of rank $r$ and type $\tau$ in $\mathcal{M}$ consists of a union of connected components $\mathcal{Z}^{+}\subset{\rm Map}(\Theta_{\tau},\mathcal{M})$, where ${\rm Map}(-)$ denotes the mapping stack (cf., [AHR20, AHR19]), such that the natural evaluation morphism ${\rm ev}_{(1,\cdots,1)}\colon\mathcal{Z}^{+}\to\mathcal{M}$ is a closed immersion.
Definition 2.37 ([HL14, Od24b]).
For a quotient algebraic $\mathbbm{k}$-stack by a linear algebraic $\mathbbm{k}$-group, a (finite) higher $\Theta$-stratification in $\mathcal{M}$ consists of
  1. ($\mathrm{i}$)
    a finite set of real numbers $1\in\Gamma\subset(-\infty,1]\subset\mathbb{R}$,
  2. ($\mathrm{ii}$)
    closed substacks $\mathcal{M}_{<c}$ of $\mathcal{M}$ for $c\in\Gamma$ which are monotonely enlarging i.e., $\mathcal{M}_{<c}\supset\mathcal{M}_{<c^{\prime}}$ if $c>c^{\prime}$ in $\Gamma$. Naturally, we can define substacks of $\mathcal{M}$ as $\mathcal{M}_{>c},\mathcal{M}_{=c},\mathcal{M}_{\geq c},\mathcal{M}_{\leq c}$.
  3. ($\mathrm{iii}$)
    higher $\Theta$-strata structure of type $N,\tau$ (Def 2.36) on each $\mathcal{M}_{=c}$ for each $c\in\Gamma$. We do not require $N,\tau$ to be identified for different $c$.
For Definition 2.36, we have the following Langton type theorem, for instance. Clearly, we can use it iteratively to obtain a generalization of [AHLH23, 6.12] for higher $\Theta$-stratification.
Theorem 2.38 (Higher $\Theta$-stable reduction [AHLH23, Od24b]).
For any quotient algebraic stack $\mathcal{M}=[H/G]$ with finite type scheme $H$ and linear algebraic group $G$ over $\mathbbm{k}$, consider any morphism $f\colon\Delta\to\mathcal{M}$ from $\Delta=\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]$, to $\mathcal{M}$ whose closed point $c$ maps into a closed substack $\mathcal{Z}^{+}$ corresponding higher $\Theta$-strata for the cone $\tau\ni\xi$, while the generic point maps outside $\mathcal{Z}^{+}$. We denote the restriction of $f$ to $\mathop{\mathrm{Spec}}\nolimits(K)$ as $f^{o}$.
Then, after a finite extension of $R$ and shrinking $\tau$, there is a toric morphism $e\colon U_{\tau}\to\mathbb{A}^{1}$ (we denote its complete localization $\mathop{\mathrm{Spec}}\nolimits\widehat{\mathcal{O}}_{U_{\tau},p_{\tau}}\to\mathop{\mathrm{Spec}}\nolimits(\widehat{\mathcal{O}}_{\mathbb{A}^{1},0}=\mathbbm{k}[[t]])$ also as $e$) and a modification of $f$ as
$f|_{\xi}\colon\mathop{\mathrm{Spec}}\nolimits\widehat{\mathcal{O}}_{U_{\tau},p_{\tau}}\to\mathcal{M},$
which extends $f^{o}\circ e$ and sends $p_{\tau}(\kappa)$ to a point outside $\mathcal{Z}^{+}$, but still isotrivially degenerates to a point in $\mathcal{Z}^{+}$.
The original statements in [Od24b] are somewhat stronger, more general and canonical, for some irrational element $\xi\in\tau$. We use the above theorem 2.38 both for the following construction of K-moduli of Calabi-Yau cones in the next section §3 and some other later work. See also [Od24c, Appendix] for another more scheme theoretic explanation on this higher $\Theta$-stable reduction and applications.
3. Moduli of Calabi-Yau cones
Now we consider the moduli of K-polystable Fano cones, i.e., Calabi-Yau cones, depending on the various preparations in the previuos section.
3.1. Boundedness
In $n=2$ ([HLQ23, LMS23]) and $n=3$ ([LMS23, Zhu23a]) case, the following boundedness result is proved, and is now generalized to any $n$ by [XZ24] 202020we learnt this result after the completion of our manuscript after the corresponding results and arguments in log Fano varieties case [Jia20].
Theorem 3.1 (boundedness [XZ24]).
For a fixed positive integer $n$, $n$-dimensional K-semistable $\mathbb{Q}$-Fano cone $X$ of the normalized volume at least $V$ are bounded.
Corollary 3.2.
In particular, for each $n$ and $V$, there are only finitely many choices of the (metric) Reeb vector fields with minimum normalized volume.
Recall that multi-Hilbert schemes [HM04] parametrizes affine closed schemes inside $\mathbb{A}^{N}$ with a linear action of an algebraic torus $T$ with fixed multi-Hilbert series. Using this, one can rephrase the above theorem:
Corollary 3.3.
For fixed $n,V$, $n$-dimensional K-semistable $\mathbb{Q}$-Fano cone $X$ of the normalized volume at least $V$ (for fixed $n,V$) are parametrized inside a finite union of certain projective multi-Hilbert schemes for all positive weights.
proof of Corollaries.
We write the proof for the convenience. For each $n$-dimensional K-semistable $\mathbb{Q}$-Fano cone $X=\mathop{\mathrm{Spec}}\nolimits\oplus_{\vec{m}}R_{\vec{m}}$ of the normalized volume at least $V$, take a positive vector field $\xi$ i.e., with $\langle\vec{m},\xi\rangle>0$ unless $R_{\vec{m}}=0$. By Theorem 3.1, combined with the latter half arguments of the proof of [Od24b, Lemma 3.14], we can take a uniform finitely generated regular submonoid $\Gamma_{\geq 0}\subset M$ generated by the set of the extremal integral vectors $S\subset\Gamma_{\geq 0}$ which contains all the moment monoids of $X$ and their $\{R_{\vec{m}}\}_{\vec{m}\in S}$ generate the coordinates ring $R$ of $X$. $S$ gives uniform embedding of $X\to\mathbb{A}^{N}$ and for small enough rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$ which includes $\xi$, we have strict positivity $\langle\tau,S\rangle\subset\mathbb{R}_{>0}$. This gives an isomorphism $T\simeq\mathbb{G}_{m}^{r}$ and so that the weights on $\mathbb{A}^{r}$ of each $\mathbb{G}_{m}$ are all positive. In particular, by [HS17, Theorem 1.1, Corollary 1.2], there is a projective multi-Hilbert scheme (possibly non-connected) which parametrizes all $n$-dimensional K-semistable Fano cone $X$ of the normalized volume at least $V$ and the Reeb vector field $\xi$. The finiteness claim of the Reeb vector fields follows from the characterization of the K-semistability [CS18] in terms of the minimized normalized volumes [LX18, 1.1], combined with the constructible semi-continuity of the minimized normalized volumes [BL21, Xu20] (See also related arguments later during the proof of Theorem 3.6). ∎
We consider the obtained finite union of projective multi-Hilbert schemes and take its union of components which parametrize those with normalized volume exactly $V$, and denote by $H_{V}$ in our paper (though it also depends on $n$ in general, just for simplicity). We sometimes fix $\xi$ among the finite choices. Note that the leading coefficient $a_{0}(\xi)$ of $F(\xi,t)$ is $V$ because of Theorem 2.7. Hence, in particular, for any pointed log terminal variety $x\in X$ parametrized in $H_{V}$, the normalized volume $\widehat{\operatorname{vol}}(x\in X)$ is at most $V$ from the definition.
Below, due to the finiteness of $\xi$, we can and do fix it in addition to the fixing of volume $V$. We then replace $T$ by the minimum algebraic subtorus which contains $\xi$ in its Lie algebra, if necessary and set $G$ to be is the commutator of $T\subset{\rm GL}(N)$, which is reductive. We then consider locus $H_{V}^{\rm ss}(\xi):=\{b\in H_{V}\mid 0\in X_{b}\text{ is K-semistable with respect to }T,\xi\}$ and replace $H_{V}$ by its closure. Note that naturally the reductive group $G$ preserves $H_{V}^{\rm ss}(\xi)\subset H_{V}$ (see e.g., [HM04, DS17]).
From now on, we want to prove that the locus of K-semistable Fano cones of the normalized volume $V$ and the Reeb vector field $\xi$, which we later denote as $B_{V}^{\rm ss}(\xi)$ makes sense and the corresponding quotient stack $[B_{V}^{\rm ss}(\xi)/G]$ admits a good moduli algebraic space which is proper.
3.2. Locally closedness
We consider the universal family over $H_{V}$ which we denote as $(\mathbb{A}^{N}\times H_{V})\supset\mathcal{U}_{V}\twoheadrightarrow H_{V}$ with the fixed multi-Hilbert series. Since the weights of the action of $T=\mathbb{G}_{m}^{r}\curvearrowright\mathbb{A}^{N}$ are all positive, the fibers contain the origin i.e., $\mathcal{U}_{V}\supset 0\times H_{V}$ and the action are good in the sense of [LS13] (if the fibers are normal).
We take the locus $B^{\prime}_{V}$ of $H_{V}$ which parametrizes normal and $\mathbb{Q}$-Gorenstein fibers by Kollár’s hull and husk [Kol08, Kol22]. (There will be also remarks later on the subtleties on definition of corresponding families and their logarithmic generalizations: Remarks 3.15, 3.16.) Further, inside $B^{\prime}_{V}$, we take the open locus $B_{V}$ where the vertices are log terminal, due to its openness.
By the standard fact (cf., e.g., [Kol13b, Lemma 3.1]), we see that for any algebraic subtorus $T^{\prime}\subset T$ of rank $1$, the $T^{\prime}$-quotients of the geometric fiber of $(\mathcal{U}_{V}\setminus\mathcal{H}_{V})\to\mathcal{H}_{V}\ni b$ are log $\mathbb{Q}$-Fano varieties precisely when $b\in B_{V}$, with respect to the natural boundary branch $\mathbb{Q}$-divisors. Hence, this $B_{V}$ is exactly the locus which parametrizes the Fano cones. We denote the restriction of the $G$-equivariant universal family to $B_{V}$ as $X_{V}\to B_{V}$. We denote an arbitrary geometric $\mathbbm{k}$-point $b\in B_{V}$ and the fiber as $X_{b}$ on which $T$ acts which commutes with the $G$-action.
Note that for any $b\in B_{V}$, $\widehat{\operatorname{vol}}(X_{b})$ is at most $V$ as we explained in the previous subsection. Therefore, if we apply the semicontinuity of the normalized volume [BL21] combined with [Xu20, 1.3] to the universal family over $B_{V}$ and the Noetherian arguments, it follows that $\{\widehat{\operatorname{vol}}(X_{b})\mid b\in B_{V}(\mathbbm{k})\}$ is a finite set which we denote as $V=V_{0}>V_{1}>\cdots>V_{m}>\cdots>V_{m^{\prime}}$. Correspondinglly, we have a filtration of $G$-invariant open subschemes
$B_{V}(v=V)\subset B_{V}(v\geq V_{1})\subset\cdots\subset\cdots\subset B_{V}(v\geq V_{m^{\prime}})=B_{V},$
where we mean
$B_{V}(v=V)$ :=$:=\{b\in B_{V}\mid\mathop{\widehat{\mathrm{vol}}}\nolimits(0\in X_{b})=V\},$
$B_{V}(v\geq V_{i})$ :=$:=\{b\in B_{V}\mid\mathop{\widehat{\mathrm{vol}}}\nolimits(0\in X_{b})\geq V_{i}\}.$
Clearly we have $B_{V}(v\geq V_{0})=B_{V}(v=V)$ though. We sometimes simply denote $B_{V}(v\geq V_{i})$ as $B_{V}(i)$.
Claim 3.4.
$B_{V}(0)=B_{V}\cap H_{V}^{\rm ss}(\xi)$.
Proof.
For a geometric point $b\in B_{V}$, the Reeb vector field $\xi$ gives a valuation $v_{\xi}$ of $X_{b}$ with $\mathop{\widehat{\mathrm{vol}}}\nolimits(v_{\xi},X_{b})=V$. Therefore, if $\mathop{\widehat{\mathrm{vol}}}\nolimits(0\in X_{b})=V$ i.e., $b\in B_{V}(0)$ if and only if the $\mathop{\widehat{\mathrm{vol}}}\nolimits$-minimizing valuation of $X_{b}\ni 0$ is exactly $v_{\xi}$ but it is characterized by the K-semistability of $(0\in X_{b}\curvearrowleft T,\xi)$ in the sense of [CS18, CS19], by [LX18, 1.1, 1.3] (cf., also [CS18, 6.1]). ∎
Hence, we can consider the algebraic stack of K-semistable Fano cones of fixed normalized volume $V$ and the Reeb vector field $\xi$, which we denote as
(22) $\mathcal{M}^{\rm ss}_{V}(\xi):=[B^{\rm ss}_{V}(\xi)/G].$
We also write $\mathcal{M}_{V}$ for the bigger quotient stack $[B_{V}/G]$, $\mathcal{M}_{V}(i):=[B_{V}(v\geq V_{i})/G]$ for each $i=1,2,\cdots,m^{\prime}$.
By [Xu20] (cf., also [BL21]), more precisely by its third paragraph of the proof of Theorem 1.3 and its Theorem 2.18, the $\mathop{\widehat{\mathrm{vol}}}\nolimits$-minimizing valuations is uniformly taken i.e., obtained as the restriction of the same quasi-monomial valuation of $\mathcal{O}_{0,\mathbb{A}^{N}}$, on each strata of some finite stratification $\{B_{V}((j))\}_{j}$ by some locally closed connected subsets $B_{V}((j))$ of $B_{V}$. Note that loc.cit crucially depends on the bounded complements by Birkar [Bir19] as well as an analogue of the invariance of local plurigenera (cf., [HMX13]). Theorem 3.6, which we next explain, further implies that each $B_{V}((j))$ is a connected component of some $B_{V}(i)\setminus B_{V}(i-1)$:
Claim 3.5.
On each connected component of $B_{V}(i)\setminus B_{V}(i-1)$, the $\mathop{\widehat{\mathrm{vol}}}\nolimits$-minimizing valuations of $0\in X_{b}$ for $b\in B_{V}(i)\setminus B_{V}(i-1)$ is uniformly taken.
We will explain this in details in the next section but related arguments are also in [Che24, Od24c].
3.3. Higher $\Theta$-reductivity-type theorem
Now we discuss $\Theta$-reductivity (cf., [HL14, AHLH23]) of the moduli stack of Fano cones in a somewhat generalized form, to include non-cone affine varieties for later purposes. The statement is compatible with the framework of [Od24b], as an irrational and family analogue of the theory of the Harder-Narasimhan filtration or the $\Theta$-strata [AHLH23].
Theorem 3.6 (cf., [Che24]).
Consider an arbitrary faithfully-flat affine klt morphism $\pi\colon Y\to S$ with a section $\sigma\colon S\to Y$ for a reduced algebraic $\mathbbm{k}$-scheme $S$ over $\mathbbm{k}$ of characteristic 00, suppose there is a constant $V$ such that for any geometric point $s\in\mathop{\mathrm{Spec}}\nolimits(R)$, the geometric fiber $Y_{s}\ni\sigma(s)$ satisfies $\mathop{\widehat{\mathrm{vol}}}\nolimits(\sigma(s)\in Y_{s})=V$.
Then there is an algebraic torus $T=N\otimes\mathbb{G}_{m}$, $\xi\in N_{\mathbb{R}}\setminus N_{\mathbb{Q}}$ and a rational polyhedral cone $\tau\ni\xi$, all independent of $s$, such that $\pi$ extends to a faithfully flat affine klt family $\tilde{\mathcal{Y}}=[\mathcal{Y}/T]$ over the quotient algebraic $S$-stack $\Theta_{\tau}\times_{\mathbbm{k}}S=[U_{\tau}(S)/T(S)]$, such that for any $s$, $\tilde{\mathcal{Y}}$ restricts to the positive weight deformations of $Y_{s}\ni\sigma(s)$ to the K-semistable Fano cones $W_{s}$ discussed in §2.4 and [Od24b, §2].
Comments on the history A half month after we posted our first version on arXiv with the above statements, [Che24] appears with a similar claim (Main Theorem 1.1 of op.cit) as above Theorem 3.6. It assertion is the existence of ideal sequence $\{\mathfrak{a}_{\lambda}\}_{\lambda\in\mathbb{R}_{\geq 0}}$ of $\mathcal{O}_{Y}$ with ${\rm Spec}_{S}\oplus_{\lambda}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ gives a family of K-semistable Fano cones, which will be our desired $\mathcal{Y}|_{p_{\tau}\times S}$. Indeed, if we take ${\rm Spec}_{S}\oplus_{\lambda\in S\subset\mathbb{R}_{\geq 0}}\mathfrak{a}_{\lambda}$ for $S:=\{\lambda\in\mathbb{R}_{\geq 0}\mid\mathfrak{a}_{\lambda}\neq\mathfrak{a}_{>\lambda}\}$, we easily see by the Nakayama’s lemma and [Tei03] (or the arguments using universal Gröbner basis cf., [Od24b]) that it gives our desired $\mathcal{Y}$ of above Theorem 3.6. Then, its author also pointed out an inaccuracy in the last step of our original proof of it (i.e., the proof of log terminality of the obtained special fibers after the proof of the necessary finite generation). Hence, we simply give up credit on this Theorem 3.6 and refer to [Che24] for the detailed proof. I thank him for the communication. His proof looks different from ours and depends on the construction of family of quasi-divisorially-log-terminal anticanonical models (“Kollár models”) on $\mathcal{Y}$ that encodes the volume-minimizing valuations (after [XZ22]). We only keep another more elementary proof for 2-dimensional case for now. (Other arguments can be also found in version 1 of this paper on arXiv, where until the middle of Step (ii) should be still valid.)
Simpler proof of Theorem 3.6 for $n=2$.
We first note that $n=2$ case can be checked by the following standard arguments, under the assumption that $S$ is a smooth $\mathbbm{k}$-curve. In this case, we take a closed point $c$ and suppose that $\sigma(c)$ has $\mathbb{Q}$-Gorenstein index $N$. Now we take the canonical cover of $Y$ as $\tilde{Y}\to Y$ of the Galois group $\mu_{N}(\overline{\mathbbm{k}})$. Suppose that the generic fiber $Y_{\eta}\ni\sigma(\eta)$ (resp., special fiber $Y_{c}\ni\sigma(c)$ for $\pi$ is the (quasi-étale) quotient by a finite group $G_{\eta}$ (resp., $G_{c}$) of the order $a_{\eta}$ (resp., $a_{c}$). If the preimage of $\sigma(S)$ in $\tilde{Y}$, denoted as $\tilde{S}$, has degree bigger than $1$ over $S$, we make base change of $\tilde{Y}\to S$ by $\tilde{S}\to S$ and denote the obtained family (resp., section) as $\tilde{\pi}\colon\tilde{Y}\to\tilde{S}$ (resp., $\tilde{\sigma}$). Then consider the pointed $\tilde{\pi}$-generic fiber which we suppose to be the quotient by a finite group $\tilde{G}_{\eta}$ (resp., $\tilde{G}_{c}$) of the order $b_{\eta}$ (resp., $b_{c}$). Note that $a_{c}\geq a_{\eta}$, $b_{c}\geq b_{\eta}$, $a_{c}=Nb_{c}$, $a_{\eta}\leq Nb_{\eta}$ from the construction. From our assumption, we have $a_{c}=a_{\eta}=\frac{4}{V}$ i.e., the local volume does not decrease at $c$. Combining them, we obtain $b_{0}=b_{\eta}$ i.e., we can assume that $\tilde{\pi}$ is a flat family of ADE singularities whose local fundamental groups orders do not change. Then it is a classical fact $\tilde{\pi}$ is formally trivial so that the (divisorial) $\mathop{\widehat{\mathrm{vol}}}\nolimits$-minimizing valuations $v_{s}$ of $\sigma(s)\in Y_{s}=\pi^{-1}(s)$ for closed points of $s\in S$ do not jump in the sense that it can be realized as a $S$-flat coherent ideal on $Y$ supported on $\sigma(S)$ (whose blow up gives a plt blow up for each $s\in S$, which corresponds to $v_{s}$). Hence, Theorem 3.6 for $n=2$ follows. ∎
In the current paper, we only need the following statements, which restricts out attention to the case of Fano cones.
Proposition 3.7.
The moduli stack $\mathcal{M}^{\rm ss}_{V}$ of $n$-dimensional K-semistable Fano cones is $\Theta$-reductive in the sense of [HL14].
proof of Proposition 3.7.
The construction of the moduli stack $\mathcal{M}^{\rm ss}_{V}$ is done in §3 (more precisely, §3.1, §3.2). Take any family of K-semistable Fano cones $\mathcal{Y}^{o}$ which comes from $(\Theta_{R}\setminus p)\to\mathcal{M}^{\rm ss}_{V}$, where $\Theta_{R}:=[\mathop{\mathrm{Spec}}\nolimits R[t]/\mathbb{G}_{m}]$ and $p$ means the closed point for the maximal ideal $(\mathfrak{m},t)$. We first restrict $\mathcal{Y}^{o}$ to the generic fiber and then denote the corresponding $\mathbb{G}_{m}$-action on the central fiber $\mathcal{Y}|_{(t)}$ as $\eta$. Then, as in [LWX21] etc, we twist the $\mathbb{G}_{m}$-action on the test configuration (without changing the total space), so that $\eta$ is replaced by $\eta+m\xi^{\prime}$ with some positive vector field $\xi^{\prime}$ and divisible enough $m$, which is still positive vector field on the fiber of $\mathcal{Y}^{o}$ over $(t)$. Then we obtain a sequence of divisorial valuations $\{E_{q}\}_{q=1,2,\cdots}$ on the generic fiber of $\mathcal{Y}^{o}$ with the center inside it as in [BX19, §2.5.1], [ABHLX20, p1026], which can be extracted by a birational morphisms $E_{q}\subset Y_{q}\to Y$ for $q\gg 0$, where $Y$ is the fiber of $\mathcal{Y}^{o}$ over $(t\neq 0)\simeq{\rm Spec}(R)$ by using [BCHM10, Blu18]. Now, we apply the same arguments as [ABHLX20, p1028-1029] to obtain an affine faithfully-flat family $\mathcal{Y}\to\Theta_{R}$ which extends $\mathcal{Y}^{o}$. As in loc.cit, we know that $(Y_{q},E_{q}+Y_{\kappa})$ is log canonical, using the ACC of log canonical threshold [HMX13]. It also follows from Lemma 2.35 applied at large enough $q$, we know the test configuration $\mathcal{Y}|_{c}$ is $\mathbb{Q}$-Gorenstein family with at worst semi-log-canonical central fiber. The remained step of the proof of Proposition 3.7 is to show that the fiber over the closed point $p=(\mathfrak{m},t)$ is a K-semistable klt cone as a special degeneration in the sense of [LX14, LWX21]. Further, note that the Donaldson-Futaki invariant (Definition 2.15 ($\mathrm{ii}$)) of the affine test configuration along $c\in S=\mathop{\mathrm{Spec}}\nolimits(R)$ is 00 as it is the same as that along the generic point $\eta\in S=\mathop{\mathrm{Spec}}\nolimits(R)$. Thus it follows from [LWX21, §4, Proposition 4.3, Corollary 4.4] that the fiber over the closed point $p=(\mathfrak{m},t)$ in $\mathcal{Y}$ is log terminal. The only remained part is to show the K-semistability of the fiber over the closed point $p=(\mathfrak{m},t)$. This follows the same arguments as [LWX21, Lemma 3.1] (a special case of the natural cone version of the CM minimization conjecture), once we modify the proof therein verbatim by replacing the Hilbert scheme by the multi-Hilbert scheme and the Futaki invariant for the Fano varieties (corresponding to regular case) by the Futaki invariant for the fixed positive vector field $\xi$. This completes the proof of Proposition 3.7.
3.4. S-completeness and its consequences
In this subsection, we confirm another ingredient for the properties of the moduli stack $\mathcal{M}_{V}^{\rm ss}(\xi)$, after [LWX21].
First we review the following algebraic stack, which is convenient for the framework of [HL14, AHLH23].
Definition 3.8 ([HL14, §2B]).
For a DVR $R$ with its uniformizer $\pi$, ${\overline{\rm ST}}_{R}:=[\mathop{\mathrm{Spec}}\nolimits(R[x,y]/(xy-\pi))/\mathbb{G}_{m}]$. Here, $\mathbb{G}_{m}$ acts on $x,y$ with weights $1,-1$ respectively. The closed point as the image of $(x,y)$ is denoted as 00.
Then, we rephrase a theorem of [LWX21] as follows.
Theorem 3.9 (cf., [LWX21]).
For any fixed $n,V$ and $\xi$, $\mathcal{M}_{V}^{\rm ss}(\xi)$ is S-complete (over $\mathbbm{k}$) in the sense of [AHLH23, §3.5, 3.38]. That is for any morphism $\varphi^{o}\colon({\overline{\rm ST}}_{R}\setminus 0)\to\mathcal{M}_{V}^{\rm ss}(\xi)$ with essentially finite type DVR $R$ over $\mathbbm{k}$, it extends to $\varphi\colon{\overline{\rm ST}}_{R}\to\mathcal{M}_{V}^{\rm ss}(\xi)$.
Proof.
This is essentially proved in [LWX21] (without the name of S-completeness in [AHLH23]). Take a $\mathbb{G}_{m}$-equivariant locally stable family of K-semistable Fano cone $\mathcal{X}^{o}\to\mathop{\mathrm{Spec}}\nolimits(R[x,y]/(xy-\pi))$ which correpsponds to $\varphi^{o}$. This is equivalent to consider two special test configurations of its general fiber $X$. Then, [LWX21, proof of 4.1] shows that it extends to (automatically $\mathbb{G}_{m}$-equivariat) faithfully flat affine $\mathbb{Q}$-Gorenstein family $\mathcal{X}\to\mathop{\mathrm{Spec}}\nolimits(R[x,y]/(xy-\pi))$. Further, [LWX21, 4.3, 4.4] (after its 4.1) shows that the fiber over the closed point $(x,y)$ is K-semistable Fano cone, which gives rise to the desired extended morphism $\varphi$. ∎
See also [BX19, ABHLX20, LX18, XZ21] for related work.
Corollary 3.10.
For a K-polystable Fano cone $T\curvearrowright X,\xi$, its automorphism group ${\rm Aut}(T\curvearrowright X,\xi)(\mathbbm{k}):=\{T\text{-equivariant automorphism }X\to X\}$ forms a reductive algebraic $\mathbbm{k}$-group ${\rm Aut}(T\curvearrowright X,\xi)$.
Proof.
This follows from the above S-completeness theorem 3.9 by [AHLH23, 3.47]. ∎
Remark 3.11.
The reductivity also follows from more differential geometric arguments, simply combining [DS17, Appendix] and [Li21, Theorem 2.9].
In §3.6, we also use the above S-completeness to prove the separatedness of the moduli, following [AHLH23, 1.1].
3.5. Properness
We are now ready to prove the following theorem, which partially generalize the results of [BHLLX21, LXZ22].
Theorem 3.12.
For any fixed $n,V$ and $\xi$, $\mathcal{M}_{V}^{\rm ss}(\xi)$ is universally closed (i.e., satisfies the existence part of the valuative criterion).
Proof.
As a direct application of Theorem 3.6, we confirm that
Claim 3.13.
the filtration $\{\mathcal{M}_{V}(i)\}_{i}$ of §3.2 naturally holds the structure of the higher $\Theta$-stratification (in the sense of [Od24b, §3.3], after [HL14]) for some rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$ which contains $\xi$.
proof of Claim 3.13.
We proceed to the proof of Claim 3.13. Take a point $b\in B_{V}(i)\setminus B_{V}(i-1)$ for $1\leq i\leq m^{\prime}$ and consider the corresponding Fano cone $X_{b}$ and its Reeb vector field $\xi_{b}\in N_{b}\otimes\mathbb{R}$ where $X_{b}$ is acted by the a priori different algebraic subtorus $N_{b}\otimes\mathbb{G}_{m}=T_{b}$ of $G$, which commutes with $T$. Corresponding to the generalization of the extended Rees construction by Teissier [Tei03, §2.1, Proposition 2.3], we consider the $T_{b}$-equivariant morphism $\psi_{b}\colon\Delta_{\xi}^{\rm gl}\to\overline{T_{b}\cdot b}$ and induced $\overline{\psi_{b}}\colon[\Delta_{\xi_{b}}^{\rm gl}/T_{b}]\to[\overline{T_{b}\cdot b}/T_{b}]$ (see also [LX18, §2.1], [Od24b, Example 2.10, Theorem 2.12]). Now, take a connected component $B$ of $(B_{V}(i)\setminus B_{V}(i-1))_{\rm red}$ which means the closed subset $(B_{V}(i)\setminus B_{V}(i-1))$ with the reduced scheme structure, regarded as a closed subscheme. Then, $N_{b}$, $T_{b}$ and $\xi_{b}$ only depends on $B$ by [Che24], so that we denote them by $N_{B},T_{B},\xi_{B}$ respectively. Recall [Od24b, Lemma 3.14] which in particular says that any element of
(23) ${\rm Map}([\Delta_{\xi_{B}}^{\rm gl}/T_{B}],\mathcal{M}_{V}(i))(B)$
[…]
comes from ${\rm Map}([\Theta_{\tau_{B}}/T],\mathcal{M}_{V}(i))(B)\text{ for some }\tau_{B}$. By Claim 3.5, there is an element of the left hand side (23) which induces degenerations to K-semistable Fano cones for any $b\in B$ simultaneously. Then, we use the above equality and represent by ${\rm Map}([\Theta_{\tau_{B}}/T],\mathcal{M}_{V}(i))(B)$ of the right hand side (24) for some $\tau_{B}$. We fix such $\tau_{B}$. Therefore, we obtain a $T_{B}$-equivariant morphism $\psi_{B}\colon U_{\tau_{B}}\times B\to B$ which induces $\overline{\psi}\colon\Theta_{\tau_{B}}\times B\to\mathcal{M}_{V}(i)$, where $U_{\tau_{B}}$ means the affine toric variety for $\tau_{B}$. $\psi$ localises to $(\Delta_{\xi}^{\rm gl}\times B)\to B_{i}$, which we still write as $\psi$. Take the connected component of the (finite type) algebraic stack ${\rm Map}([\Theta_{\tau_{B}}/T],\mathcal{M}_{V}(i))$ (cf., [HL14], [AHR20, 5.10, 5.11], [AHR19, 6.23]) which contains the image of $B$ and denote as $\mathcal{B}$. Take any geometric $\mathbbm{k}$-point $b^{\prime}\in\mathcal{B}$ and consider the corresponding image of the vertex $p_{\tau_{B}}$ of $U_{\tau_{B}}$ as $[Y\subset\mathbb{A}^{N}]$. Then from the construction, its multi-Hilbert function remains the same as those parametrized by $B$. So, combined with the uniqueness of the $\mathop{\widehat{\mathrm{vol}}}\nolimits$-minimizer [XZ21, BLQ22], it follows that the images of $p_{\tau_{B}}\times B$ are again inside $B_{i}\setminus B_{i-1}$ and the image of $\Theta_{\tau}\times\mathcal{B}\to\mathcal{M}_{V}$, which extends $\overline{\psi}$, still lies inside $\mathcal{M}_{V}(i)$. Therefore, we obtain the higher $\Theta_{\tau_{m}}$-stratification structure on each strata on $\mathcal{M}_{V}$. ∎
Now, we proceed to the proof of Theorem 3.12 by the valuative criterion. Take any DVR $R$ of essentially finite type over $\mathbbm{k}$ (cf., [AHLH23, Appendix A3]), and denote its quotient field (resp., residue field) as $K$ (resp., $\kappa$) and set $S:={\rm Spec}(R)$, its generic point $\eta$ (resp., closed point $c$). Consider an arbitrary morphism $\varphi\colon S\to B_{V}$ which maps $\eta$ into $B_{V}^{\rm ss}(\xi)$. We denote the corresponding family of Fano cones as $Y_{K}\to\eta$. Now we want to show that the induced morphism $\eta\to\mathcal{M}^{\rm ss}_{V}(\xi)$ can be extended to a morphism $S\to\mathcal{M}^{\rm ss}_{V}(\xi)$. Take any common rational positive vector field $\xi^{\prime}$ and corresponding algebraic subtorus $T^{\prime}\subset T$ of rank $1$. Now, we apply [LX14, Theorem 1] $T$-equivariantly to a $\mathbb{Q}$-Gorenstein family of log Fano pairs $Y_{K}/T^{\prime}\twoheadrightarrow\eta$ (cf., which underlies the associated Seifert $\mathbb{G}_{m}$-bundle’s base cf., [Kol13b]) and take its cone: possibly after replacement of $R$ by its finite extension, this gives an extension $T\curvearrowright Y\twoheadrightarrow S$ of $(T/T^{\prime})\curvearrowright Y_{K}/T^{\prime}\twoheadrightarrow\eta$ as a $T$-equivariant family of Fano cones, corresponding to some morphism $S\to B_{V}(i)$ (after temporarily enlarging $B_{V}$ if necessary). Now, by Claim 3.13, we can use the higher $\Theta$-stable reduction theorem in [Od24b, §3.2, §3.3], in the place of [AHLH23] for the original proof of the Fano varieties case [BHLLX21, LXZ22] to decrease $i$, and hence finally reach $i=0$ by its repetition thanks to the ACC of normalized volumes [XZ24, 1.2]. This completes the proof of Theorem 3.12. ∎
3.6. Existence of good moduli space
3.6.1. General statements and proof
Now we are ready to construct the moduli spaces.
Theorem 3.14 (Proper K-moduli of Fano cone).
For fixed $n,V$, there is a moduli Artin stack of $n$-dimensional K-semistable Fano cones of normalized volume $V$ and $\mathcal{M}_{V}^{\rm ss}(\xi)$, which admit a proper good moduli space $M_{V}^{\rm ss}(\xi)$ whose closed points parametrize K-polystable Fano cones among them.
Proof.
We constructed the moduli stack of K-semistable Fano cones of normalized volume $V$ and $\mathcal{M}_{V}^{\rm ss}(\xi)$ as (22). Then by the $\Theta$-reductivity (Proposition 3.7 of Theorem 3.6) and the S-completeness theorem 3.9 (cf., [LWX21, §4]) of $\mathcal{M}_{V}^{\rm ss}(\xi)$, [AHLH23, Theorem A] can be applied to prove the existence of separated good moduli space $M_{V}^{\rm ss}(\xi)$.
Furthermore, recall that the normalized volumes of $n$-dimensional log terminal singularities satisfy ACC [XZ24, 1.2]. Thus, Theorem 3.12 as an application of the higher $\Theta$-stable reduction theorem [Od24b, §3] can be applied to show that $M_{V}^{\rm ss}(\xi)$ is universally closed and hence proper. We complete the proof. ∎
Remark 3.15 (Moduli ($2$-)functor).
The set-theoretic meaning of the moduli $\mathcal{M}_{V}^{\rm ss}(\xi)(\mathbbm{k})$ and $M_{V}^{\rm ss}(\xi)(\mathbbm{k})$ are clear i.e., $n$-dimensional K-semistable (resp., K-polystable) Fano cones of volume density $V$ and the Reeb vector field $\xi$ for $T$, as we defined (in the previous section), but more generally $S$-valued points for more general $\mathbbm{k}$-schemes $S$ i.e., $\mathcal{M}_{V}^{\rm ss}(\xi)(S)$ can be redefined more intrinsically as their “locally stable” families $T\curvearrowright\mathcal{X}\twoheadrightarrow S$ in the sense of Kollár [Kol22, Definition 3.40]. The arguments are essentially the same as [Kol08, Kol22, Xu20], so we omit the detail.
Remark 3.16 (Generalization to logarithmic setup).
[…]
Then, by the log boundedness of $\mathcal{M}_{V}^{\rm lss,red}(\xi,l)(\mathbbm{k})$ ([XZ24]), we use the multi-Hilbert scheme [HS17] again of $T\curvearrowright X$ together with [Kol22, 7.3] (applied to projective compactifications of $D$s with respect to a positive vector field $\xi^{\prime}\in N$) to obtain a larger moduli stack of log affine cones, in the sense of Definition 2.9. Note that the divisors $D$ are also assumed to be $T$-invariant which, again by the theory of [HS17], provides the reason of locally finite typeness of the obtained stack, and its finite typeness is ensured by the log boundedness [XZ24]. Then, by [Kol22, 3.22] (also cf., [Kol08]), we can take its locally closed substack generalize our previous construction of $\mathcal{M}_{V}^{\rm ss}(\xi)$. Then, the remained confirmation of its properties work verbatim after our discussions above in this §3 (and the materials of §2 and the references used in the proofs), using the logarithmic framework of modern birational geometry (cf., [KM98, §3]). Summarizing up, our arguments in this paper gives the following, similarly to Theorem 3.14:
Theorem 3.17 (K-moduli of log Calabi-Yau cones).
We fix $n,l\in\mathbb{Z}$, and $V\in\mathbb{R}_{>0}$. Then, by the logarithmic version of the boundedness with fixed $n,N,V,l$. ([Jia20, XZ24], cf., also Corollaries 3.2, 3.3), there are only finitely many choices of $N,T=N\otimes G_{m},\xi\in N\otimes\mathbb{R}$ of $n$-dimensional log K-semistable log Fano cones $(T\curvearrowright X\ni x,\Delta=\frac{1}{l}D)$ where $(X,\Delta)$ are kawamata-log-terminal and are of the normalized volume $V$.
Further fixing $N,T=N\otimes G_{m},\xi\in N\otimes\mathbb{R}$ (just for convenience of description), they form a moduli Artin stack $\mathcal{M}_{V}^{\rm lss,red}(\xi,l)$ at the reduced Artin stack level, which admits a reduced proper good moduli space $M_{V}^{\rm lss,red}(\xi,l)$ whose closed points parametrize $n$-dimensional log K-polystable log Fano cones $(X,\frac{1}{l}D)$ among them.
In particular, if $\mathbbm{k}=\mathbb{C}$, $M_{V}^{\rm lss,red}(\xi,l)(\mathbb{C})$ can be again regarded as the moduli of Sasaki-Einstein manifolds with certain singularities including “conical” type which we do not give intrinsic formulation in this paper.
3.6.2. Reconstructing Fano varieties’ K-moduli
For the case when $\xi\in N$ and provides regular positive (Reeb) vector field, the above theorem 3.14 reproves the following.
Corollary 3.18 (K-moduli of $\mathbb{Q}$-Fano varieties).
For fixed $n$ and fixed $V$, there is a moduli stack $\mathcal{M}$ of $n$-dimensional K-semistable $\mathbb{Q}$-Fano varieties $X$ of the anticanonical volume $V=(-K_{X})^{n}$, and it further admits a proper good moduli space $M_{V}^{\rm ss}(\xi)$ whose closed points parametrize K-polystable $\mathbb{Q}$-Fano varieties.
We make further discussions on the possibility of yet other (general) apporoaches to algebraic construction of K-moduli of Fano varieties, in the next section §4.
Proposition 3.19.
We fix $n,V,T,\xi$ below. The (universal) CM $\mathbb{R}$-line bundle $\lambda_{\rm CM}(T\curvearrowright\mathcal{X})$ (Definition 2.22) on the moduli stack $\mathcal{M}_{V}^{\rm ss}(\xi)$ of $n$-dimensional Fano cones of the volume density $V$ for the Reeb vector field $\xi$ descends to a $\mathbb{R}$-line bundle on the good moduli space $M_{V}^{\rm ss}(\xi)$ which we denotes as $L_{\rm CM}(n,V,\xi)$ for simplicity.
Proof.
The proof is essentially the same arguments as [OSS16, §6.2, after (K-moduli) Conjecture 6.2]. Indeed, since the morphism $\mathcal{M}_{V}^{\rm ss}(\xi)\to M_{V}^{\rm ss}(\xi)$ is étale locally a GIT quotient as in the discussion of [OSS16], the descendability is equivalent to the vanishing of the Donaldson-Futaki invariants for product test configurations for every K-semistable Fano cones, which follows from the definition of K-semistability ([CS18], our Definition 2.15). ∎
3.6.3. Projectivity and CM $\mathbb{R}$-line bundle
Naturally, as conjectured in polarized setup in [FS90], [Od10], [OSS16, 6.2] and algebraic parts solved by [CP21, XZ22] for Fano varieties case, we conjecture:
Conjecture 3.20.
For $\mathbbm{k}=\mathbb{C}$, $c_{1}(M_{V}^{\rm ss}(\xi)(\mathbb{C}),L_{\rm CM}(n,V,\xi))$ is represented by a Weil-Petersson type Kähler current on $M_{V}^{\rm ss}(\xi)^{\rm an}=M_{V}^{\rm ss}(\xi)(\mathbb{C})$. For general $\mathbbm{k}$, $L_{\rm CM}(n,V,\xi)$ is an ample $\mathbb{R}$-line bundle so that $M_{V}^{\rm ss}(\xi)$ is projective.
We omit the precise form of the natural log extension of the CM line bundle and the corresponding positivity conjecture, which the readers can guess from the polarized varieties case (cf., e.g., [OSS16, §6], [CP21, §3]).
3.7. Examples
3.7.1. Quasi-regular case
In the quasi-regular case, as it is naturally expectable after [MSY06, MSY08, CS18, CS19], we can reduce the study to the K-moduli of (proper) log Fano varieties. This generalizes Corollary 3.18, which corresponds to the regular Reeb vector field case.
Proposition 3.21.
The moduli stack $\mathcal{M}_{V}^{\rm ss}(\xi)$ (resp., $M_{V}^{\rm ss}(\xi)$) of $n$-dimensional Fano cones $(T=\mathbb{G}_{m})\curvearrowright X$ of volume density $V$ for quasi-regular Reeb vector field $\xi\in N_{\mathbb{Q}}$ is isomorphic to the K-moduli stack (resp., K-moduli space) of certain log Fano varieties which appears as quotient of $X$.
Proof.
Take any $\mathbb{G}_{m}$-equivariantly faithfully flat family of affine cones $\mathbb{G}_{m}\curvearrowright\mathcal{X}=\cup_{s}X_{s}\subset\mathbb{A}^{N}\otimes S\to S\ni s$ with the weight vector $\xi=(m_{1},\cdots,m_{N})$, where $m_{i}$ are coprime positive integers. By diagonalization, any $\mathbb{G}_{m}$-equivariantly faithfully flat family of quasi-regular affine cones can be localized to the above type family. As noted by [CS18, CS19] (cf., also [LL19]), the K-(semi/poly)stability of $X_{s}$ is equivalent to that of $(X_{s}\setminus x)/\mathbb{G}_{m}$ with $\sum_{D}\frac{m_{D}-1}{m_{D}}D$ where $D$ runs over prime divisors inside $(z_{i}=0)\subset(X_{s}\setminus x)/\mathbb{G}_{m}$ and $m_{D}$ is the ramification degree of $[(X_{s}\setminus x)/\mathbb{G}_{m}]\to(X_{s}\setminus x)/\mathbb{G}_{m}$ i.e., $m_{D}={\rm gcd}(m_{1},\cdots,m_{i-1},m_{i+1},\cdots,m_{N})$.
If we consider the K-moduli stack $\mathcal{M}$ of log Fano varieties $((X_{s}\setminus x)/\mathbb{G}_{m},\sum_{D}\frac{m_{D}-1}{m_{D}}D)$, the above arguments give us a morphism $\varphi\colon\mathcal{M}_{V}^{\rm ss}(\xi)\to\mathcal{M}$. The inverse of this morphism also exists by taking the relative cones. Hence, these moduli stacks are isomorphic. Further, because of the fact that the K-polystability of $X$ and $\varphi(X)$ are equivalent, or otherwise from the existence of proper good moduli spaces for both, we conclude that $\varphi$ also induces an isomorphism at the good moduli space level. ∎
Therefore, we can reduce the following to the log K-stability and log K-moduli of (usual) log Fano varieties.
Problem 3.22.
Explicitly describe the structure of K-moduli (Theorem 3.14) of quasi-regular Fano cones (Sasaki-Einstein manifolds with singularities) in the case of [BGN03, BG05, Kol05, LST22].
3.7.2. Irregular case
The examples of compact irregular Sasaki-Einstein manifolds found in the initial stage seem to be [GMSW04a], [GMSW04b], [MS21], [FOW09] among others and they are mostly toric Fano cones, so that they do not have positive dimensional moduli spaces.
Recently, Süß[Sus21] considered an example of $3$-dimensional affine $T$-varieties of complexity one (cf., [IS17]) which admit a positive dimensional moduli space. The example in loc.cit is constructed as follows.
By the correspondence of normal affine $T$-varieties with poloyhedral divisors [AH06], if we consider $Y=\mathbb{P}^{1}$ and a polyhedral divisor ([AH06]) i.e., $\mathcal{D}=\sum_{y\in\mathbb{P}^{1}(\mathbbm{k})}\mathcal{D}_{y}$ where $\mathcal{D}_{y}$ are rational polyhedra of the same tail cone $\sigma$ which satisfy “proper”ness condition of [AH06, §1].
In [Sus21, §6], disjoint $y_{1},\cdots,y_{k}\in\mathbb{P}^{1}(\mathbbm{k})\setminus\{0,\infty\}$ are fixed and we set $\sigma$, $\mathcal{D}_{y}$ as:
  • $\sigma:=\mathbb{R}_{\geq 0}(-1,1)+\mathbb{R}_{\geq 0}(15k-4,8)$,
  • $\mathcal{D}_{0}:=(\frac{2}{5},\frac{1}{5})+\sigma$,
  • $\mathcal{D}_{\infty}:=(\frac{-2}{3},\frac{1}{3})+\sigma$,
  • $\mathcal{D}_{y_{i}}(i=1,\cdots,k):=([0,1]\times\{0\})+\sigma$,
  • $\mathcal{D}_{y}:=\sigma$ otherwise i.e., when $y\neq 0,\infty,y_{i}(i=1,\cdots,k)$.
In [Sus21, §6], the corresponding $T$-variety $X_{k}$ is proven to be K-polystable Fano cone for any positive integer $k$. Due to the automorphism group $\mathbb{G}_{m}(\mathbbm{k})=\mathbbm{k}^{*}$ of $(\mathbb{P}^{1},(0)+(\infty))$, we have the corresponding moduli
(25) $M_{k}(\mathbbm{k})=(\mathbb{P}^{1}(\mathbbm{k})\setminus\{0,\infty\})^{k}/(\mathbb{G}_{m}(\mathbbm{k})=\mathbbm{k}^{*})\simeq\mathbbm{k}^{k-1}.$
To explicitly understanding our compactification (Theorem 3.14) in this case, now we would like to consider the case when $y_{i}$s can collide and can be even 00 or $\infty$. When the supports of some $y$ collide, we take the Minkowski sum of the original $\mathcal{D}_{y}$s.
Case 1
Firstly, we consider the effect of collisions of $y_{i}$ i.e., $y_{i}=y_{j}$ outside $0,\infty$ for some $1\leq i\neq j\leq k$.
By [IS17], [Sus21, Definition 4.2, Proposition 4.3], special test configuration of the obtained $T$-variety $X$ corresponds to an admissible point $y\in\mathbb{P}^{1}$ in the sense of loc.cit. Further, by [Sus21, Theorem 4.10], the K-polystability of $(X,\xi)$ remains equivalent as far as $\sigma_{y}$ in loc.cit §4 and ${\mathbf{u}_{y}}$ does not change. Thus, in particular, $X$ is K-polystable if $y_{i}$s are all in $\mathbb{P}^{1}\setminus\{0,\infty\}$ even if they collides.
Case 2
Next we consider when some of $y_{i}$ collide with 00. Suppose exactly $m(0\leq m\leq k)$ of $y_{i}$s attain 00. Then, $\mathcal{D}_{0}$ is changed to $(\frac{2}{5},\frac{1}{5})+([0,m],0)\sigma$.
Consider the case with $k=2$ and positive $m$. A lengthy hand calculation (after [Sus21, 4.10, §6]), which we omit here, and confirmation by using the Sage package “TVars” created by Absar Gull, Leif Jacob, Leandro Meier and Hendrik Süß, shows the Donaldson-Futaki invariant of the special test configuration for the admissible $y=0$ is negative. The author thanks the creators of the Sage package, especially Hendrik Süßfor the help on this computation.
Case 3
Next we consider when some of $y_{i}$ collide with $\infty$. Then, if $k=2$ and $m^{\prime}$ is positive, a similar calculation as above (after [Sus21, 4.10, §6] and the same Sage program) shows the Donaldson-Futaki invariant of the special test configuration for the admissible $y=\infty$ is positive if $m^{\prime}=1$ and negative if $m^{\prime}=2$.
The conclusion is that
Proposition 3.23.
In the generalized setup of [Sus21, §6] (allowing $y_{i}$ to be $0,\infty$ and their collision), suppose $k=2$ and $y_{i}=0$ for $m$ $i$s and $y_{i}=\infty$ for $m^{\prime}$ $i$s. Then, the obtained Fano cone $X=X_{k}$ is K-polystable if and only if $m\leq 1$.
Corollary 3.24.
For $k=2$, the normalization of the K-moduli compactification $M_{k}=M_{2}$ (Theorem 3.14) of the above moduli 3.23 ((25)) of the $3$-dimensional $T$-variety Fano cone of complexity $1$ is $\mathbb{P}^{1}=\mathbb{P}(1,2)$.
proof of Corollary 3.24.
We consider $\frac{1}{y_{i}}=:x_{i}$ and take $[x_{1}+x_{2}:x_{1}x_{2}]\in\mathbb{P}(1,2)=(\mathbb{A}^{2}_{x_{1}+x_{2},x_{1}x_{2}}\setminus(0,0))/\mathbb{G}_{m}$. The natural relative version of the construction of $T$- variety in [AH06] gives a family of $T$-variety $X$ over the above $\mathbb{A}^{2}_{x_{1}+x_{2},x_{1}x_{2}}\setminus(0,0)=(\mathbb{A}^{2}_{x_{1},x_{2}}\setminus(0,0))/S_{2}$. By combining with Proposition 3.23, we obtain the proof. ∎
By similar computations and some experiments, we expect that for at least $k=3,4$ and possibly other small $k$s, that the normalization of the K-moduli compactification of $M_{k}$ is isomorphic to the log K-moduli of
$(\mathbb{P}^{1},(1-c)[0]+c[y_{1}]+\cdots+c[y_{k}]+(1-3c)[\infty])$
for distinct $y_{i}$s and $0<c\ll 1$ and their log K-polystable degenerations.
It would be also interesting to work on these moduli from quiver-gauge theoretic side on the ADS-CFT correspondence.
4. Discussions
4.1. Via $\gamma$-invariants of Berman
The reason of the name of “$\delta$”-invariant in [FO18] comes from the original work of Berman [Ber13], where he introduces “$\gamma$”-invariant $\gamma(X)\in\mathbb{R}_{>0}$ for $\mathbb{Q}$-Fano varieties $X$ from his original probabilistic (or statistical mechanical) approach to the complex Monge-Ampère measures of the Kähler-Einstein metrics. See [FO18, §2.2] for the more algebraic explanation and comparision of the two invariants ($\gamma$ vs $\delta$). We also recall that
Theorem 4.1 ([FO18, Theorem 2.5]).
$\delta(X)\geq\gamma(X)$ for any $\mathbb{Q}$-Fano variety $X$.
See [Ber13] for the discussions on the analytic aspects of the possibility that the above can be equality in general. We point out below that the theory of $\gamma(X)$ a priori gives a direct algebraic construction of K-stable limit of Fano varieties without the iterative procedure [AHLH23, Od24b] nor some invariants a priori, as in the method via $\delta(X)$.
Proposition 4.2 (K-stable Fano via log KSBA theory).
[…]
Recall that standard birational geometric arguments easily imply the above construction does not depend on the choice of $f$. Thus, the main point of the above observation is its construction of K-(semi)stable filling $\mathcal{X}_{0}$ does not depend on the choice of divisors etc, nor any kind of iterative procedures, since $\mathcal{D}_{m}$ is taken as a canonical object (for each $m$). Hence, this approach is close to that of the Kollár-Shepherd-Barron-Alexeev for the (log) K-ample case (cf., [Kol22]), which we now can interpret as a part of (log) K-stability theory (cf., the old survey [Od10] and references therein).
4.2. Via $\delta$-invariants
Recall that the original proof in [BHLLX21, LXZ22] crucially uses the following refinement of $\delta$-invariant:
Definition 4.3 ([BHLLX21, (1.2)]).
For each log terminal $\mathbb{Q}$-Fano variety $X$,
$M^{\mu}(X):=\biggl{(}\delta(X),\quad\inf_{\mathcal{X}}\dfrac{\rm DF(\mathcal{X},-K_{\mathcal{X}/\mathbb{P}^{1}})}{||\mathcal{X}||_{L^{2}}}\biggr{)},$
as an element of $\mathbb{R}_{\geq 0}^{2}$ to which we put the lexicographic order. Here, $\delta(X)$ denotes the $\delta$-invariant of $(X,-K_{X})$ ([FO18, BJ22]), $\mathcal{X}$ runs over special test configurations of $X$, $||\mathcal{X}||_{L^{2}}$ is the $L^{2}$-norm of test configuration $(\mathcal{X},-K_{\mathcal{X}/\mathbb{P}^{1}})$.
The crux of the original algebraic proof of this theorem is that we can modify the family $\mathcal{X}$ or its corresponding family $f\colon\Delta\to\mathcal{M}$ to improve $M^{\mu}(f(0))$ to make the first component $\delta(f(0))$ at least $1$.
Comparison of the $\delta$-invariant ([FO18, BJ22]) and the normalized volumes ([Li18, Li17, LL19]) are rather nontrivial problems, although they are analogous ([Liu18], [LLX18, Remark 4.5]). For instance, the following is known.
Theorem 4.4 (cf., [Liu18],[BBJ21],[CRZ19],[BJ22]).
For any $\mathbb{Q}$-Fano variety $X$, we have
(26) $\delta(X,-K_{X})^{n}(-K_{X})^{n}\leq\biggl{(}1+\frac{1}{n}\biggr{)}^{n}\mathop{\widehat{\mathrm{vol}}}\nolimits(x\in X).$
A weaker version is that
(27) $\min\{\delta(X,-K_{X}),1\}^{n}(-K_{X})^{n}\leq\biggl{(}1+\frac{1}{n}\biggr{)}^{n}\mathop{\widehat{\mathrm{vol}}}\nolimits(x\in X).$
Proof.
The weaker version (27) follows from the combination of the fact that $\min\{\delta(X),1\}$ is the greatest Ricci lower bound ([BBJ21, §7.3], [CRZ19, Appendix]) and the logarithmic analogue [LL19, Proposition 4.6] of [Liu18, Theorem 2], applied to $(X,(1-\min\{\delta(X,-K_{X}),1\})D)$. The stronger version (26) also directly follows from [BJ22, Theorem D] if we set $L=-K_{X}$ for $X$. ∎
On the other hand, recently the theory of the $\delta$-invariants are also generalized to the (possibly irregular) Fano cone setups in a similar manner by [Wu22, Hu23]. Note that [Hu23, 7.0.1] and the fact that minimized normalized volume can be irrational would imply that they are different in general. A natural question is
Question 4.5.
Can we use the above-mentioned generalizations of $\delta$-invariant ([Wu22, Hu23]) to also give another properness proof of K-moduli of Fano cones, extending the approach in [BHLLX21, LXZ22]?
We plan to discuss versions of our results for Kähler-Ricci solitons in a different paper, which can be often considered as special cases of the Sasaki-Einstein manifolds (cf., [MN15, Conjecture 1.2]).
Acknowledgement
The author thanks R.Berman, S.Sun, H.Süß, and C.Spotti for the helpful conversations and inspirations related to this paper. He also thanks Z.Chen, M.Jonsson, M.Yamazaki for comments on the manuscript. During the research, the author was partially supported by Grant-in-Aid for Scientific Research (B) 23H01069, 21H00973, Grant-in-Aid for Scientific Research (A) 21H04429, 20H00112, and Fund for the Promotion of Joint International Research (Fostering Joint International Research) 23KK0249. The last part of the work is done during the author’s stay at Institut de mathématiques de Jussieu, Paris Rive Gauche in Paris. We appreciate all of them who helped our pleasant stay for the kind hospitality.